### v1

# AsymDex: Leveraging Asymmetry and Relative Motion in Learning Bimanual Dexterity

Abstract
We present Asymmetric Dexterity (AsymDex), a novel reinforcement learning (RL) framework that can efficiently learn asymmetric bimanual skills for multi-fingered hands without relying on demonstrations, which can be cumbersome to collect. Two crucial ingredients enable AsymDex to reduce the observation and action space dimensions and improve sample efficiency. First, AsymDex leverages the natural asymmetry found in human bimanual manipulation and assigns specific and interdependent roles to each hand: a facilitating hand that moves and reorients the object, and a dominant hand that performs complex manipulations on said object. Second, AsymDex defines and operates over relative observation and action spaces, facilitating responsive coordination between the two hands. Further, AsymDex can be easily integrated with recent advances in grasp learning to handle both the object acquisition phase and the interaction phase of bimanual dexterity. Unlike existing RL-based methods for bimanual dexterity, which are tailored to a specific task, AsymDex can be used to learn a wide variety of bimanual tasks that exhibit asymmetry. Detailed experiments on four simulated asymmetric bimanual dexterous manipulation tasks reveal that AsymDex consistently outperforms strong baselines that challenge its design choices, in terms of success rate and sample efficiency. The project website is at https://sites.google.com/view/asymdex-2024/.
Keywords: Dexterous Manipulation, Bimanual Manipulation
1 Introduction
We tackle the challenge of learning bimanual dexterous manipulation skills on multi-fingered hands using reinforcement learning (RL). Bimanual skills are crucial for robots operating in human environments as they allow for more complex and flexible manipulation compared to a single hand [1, 2, 3, 4, 5, 6, 7]. While learning dexterous manipulation skills on a single hand presents numerous challenges [8, 9, 10, 11, 12, 13, 14], learning bimanual dexterity can be significantly more challenging due to the higher-dimensional state and action spaces and the need to coordinate and synchronize the movement of two hands [1].
To circumvent the challenges introduced by the high-dimensional state and action spaces of bimanual dexterous manipulation, we take inspiration from how humans approach this challenge. Humans exhibit a natural asymmetry in how we use each of our hands when we perform most bimanual tasks. Specifically, we tend to use one hand to reposition and reorient an object being manipulated so as to make it easier for the other hand to achieve the desired manipulation objectives. While the asymmetric assumption might seem restrictive at first glance, rich bodies of work in human biomechanics and evolution reveal its significance and necessity [15, 16, 17, 18]. Evolutionary biologists posit that such handedness evolved in humans and great apes to meet the escalating cognitive demands of tool use and complex manipulation [19]. Indeed, a large class of real-world bimanual tasks admit this asymmetry (e.g., attachment, detachment, assembly, pouring).
A key insight that we leverage is that the natural asymmetry between the two hands can reduce the dimensionality of the bimanual dexterous manipulation, and in turn improve effectiveness and sample efficiency. Using this insight, we contribute a novel learning framework, dubbed Asymmetric Dexterity (AsymDex), that can efficiently learn asymmetric bimanual dexterous manipulation tasks based on RL (see Fig. 1 for a block diagram). Note that AysmDex does not require the cumbersome collection of demonstrations and can learn bimanual skills only using reinforcement learning.
Figure 1: Our approach (AsymDex) efficiently learns asymmetric bimanual dexterous manipulation skills based on reinforcement learning by effectively leveraging i) the natural asymmetry in the hands’ roles and ii) relative state and action spaces that encourage synchronization.
AsymDex has two crucial ingredients. First, it introduces asymmetry by defining a dominant hand and a facilitating hand. While the facilitating hand learns to reposition and reorient the object, the dominant hand learns complex manipulation skills (including in-hand manipulation). Note that there’s no relative movement between the grasped object and the facilitating hand. As a result, AsymDex holds the fingers of the facilitating hands in the grasping pose and only controls the 6D motion of its base. On the other hand, AsymDex learns to control both the base and fingers of the dominant hand. Second, AsymDex reasons about and controls the relative motion between the dominant and facilitating hands. We define relative observation and action spaces that incentivize responsive coordination between the two hands without resorting to explicit time-dependence.
We also leverage the observation that bimanual manipulation in practice is composed of two distinct phases: i) the acquisition phase in which objects are grasped from surfaces, and ii) the interaction phase where the two hands coordinate to perform the bimanual task. Unlike many existing methods that ignore the acquisition phase, we show that this decomposition enables AysmDex to be seamlessly integrated with learned grasping policies to enable fluent execution.
In summary, we contribute AsymDex – a novel framework for asymmetric bimanual dexterity that leverages the natural asymmetry in hand roles and relative observation and action spaces. We evaluate AsymDex on four complex bimanual dexterous tasks (adapted from BiDexHand [20]) and compare against strong baselines that challenge its design choices. Our results show that AsymDex consistently outperforms the baselines in terms of success rate and sample efficiency.
2 Related Work
In this section, we contextualize our contributions within relevant sub-fields.
Learning Bimanual Manipulation: Several existing methods focus on learning bimanual skills, but are often limited to simple end-effectors. Imitation learning (IL) based approaches have been particularly successful in bimanual manipulation [21, 22, 3, 23], and have led to novel and low-cost infrastructure to collect bimanual manipulation data [4, 24]. These approaches rely on demonstrations to provide the necessary supervision to learn effective coordination strategies. Reinforcement learning (RL) has also been shown to be successful in learning bimanual manipulation skills [25, 26, 27, 28]. These methods implicitly incentivize coordination by learning to optimize reward functions that favor task success and efficiency. In contrast to all of these works that only consider parallel jaw grippers, AsymDex learns bimanual dexterous manipulation skills involving multi-fingered hands.
Asymmetry in Bimanual Manipulation: Motivated by the asymmetry in how humans use their two hands (referred to as role-differentiated bimanual manipulation [16, 17, 18, 29]), recent works assign different roles to each robot hand in the bimanual system [30, 31, 32, 2, 33]. However, some of these approaches restrict the role of the facilitating hand to stabilizing the object while the dominant hand manipulates it [30, 31, 2]. In contrast, AsymDex allows the facilitating hand to reposition and reorient the object simultaneously as the dominant hand executes its role. Importantly, unlike AsymDex, all these prior methods are limited to parallel jaw grippers.
Learning Dexterous Manipulation: Learning dexterous manipulation skills involves addressing numerous challenges due to high dimensional state and action spaces and highly nonlinear dynamics. Recent works have tackled these challenges using imitation learning (IL) or reinforcement learning (RL) and demonstrate impressive performance [8, 9, 34, 10, 35, 11, 12, 36, 37, 14]. However, IL-based methods, including those that combine RL and IL [8, 10], rely either on complex infrastructure and retargeting methods to collect demonstrations  [8, 35, 14, 38, 39] or pre-trained expert policies [11, 12, 37]. On the other hand, RL-based methods do not share these constraints as they learn skills via reinforcement. However, RL methods tend to require single amounts of exploration even for unimanual dexterous manipulation [9, 34, 10, 36]. As we show in our experiments, naive application of RL-based methods is not effective for bimanual dexterous manipulation due to the increased dimensionality and the need for coordination. We also show that AsymDex is able to efficiently learn complex bimanual dexterous manipulation skills using RL without relying on demonstrations.
Learning Bimanual Dexterous Manipulation: A few recent studies have focused on learning bimanual dexterity. Some of these methods require the collection of expert demonstrations [6] and suffer from the same limitations we discussed earlier for IL-based methods that use parallel jaw grippers. To circumvent the need for collecting demonstrations, recent efforts have led to methods that only leverage RL and yet are capable of learning impressive bimanual manipulation skills, such as playing the piano [7], twisting lids off containers [40], and dynamic handover [5]. While these methods are specifically designed to solve a particular task, AsymDex is capable of efficiently learning different bimanual dexterous manipulation tasks.
3 Learning Asymmetric Bimanual Dexterous Manipulation Skills
In this section, we formulate the problem of asymmetric bimanual dexterous manipulation and introduce the different elements of our approach (AsymDex).
3.1 Preliminaries
Consider the problem of bimanual dexterous manipulation, in which two multi-fingered hands coordinate to manipulate up to two objects. Formally, this problem can be defined as a Partially-Observable Markov Decision Process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{Z},\mathcal{A},\mathcal{R},\mathcal{P})$, where $\mathcal{S}\in\mathbb{R}^{n}$ is the state space, $\mathcal{Z}\in\mathbb{R}^{m}$ is the observation space, $\mathcal{A}\in\mathbb{R}^{u}$ is the action space, ${\mathcal{R}:\mathbb{R}^{m}\times\mathbb{R}^{u}\rightarrow\mathbb{R}}$ is the reward function, and ${\mathcal{P}:\mathbb{R}^{n}\times\mathbb{R}^{u}\rightarrow\mathbb{R}^{n}}$ is the environment dynamics. Note that we do not assume access to any demonstrations. Instead, we tackle of challenge of learning purely based on reinforcement. Given this formulation, the problem boils down to learning a policy $\pi:\mathcal{Z}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $E_{\pi}[\Sigma_{t=0}^{T-1}\gamma^{t}\mathcal{R}(z(t),a(t))]$.
Observation and Action Space: The observation space $\mathcal{Z}$ is composed of hand and object measurements. At step $t$, $z(t)$ contains $\xi^{b}_{1}(t)\in\mathbb{R}^{6}$, $\xi^{b}_{2}(t)\in\mathbb{R}^{6}$, $\xi^{h}_{1}(t)\in\mathbb{R}^{n_{1}}$, $\xi^{h}_{2}(t)\in\mathbb{R}^{n_{2}}$, $o_{1}(t)\in\mathbb{R}^{6}$, and $o_{2}(t)\in\mathbb{R}^{6}$, where $\xi^{b}_{1}$ and $\xi^{b}_{2}$ denote the first and the second hand bases’ $6D$ poses, $\xi^{h}_{1}$ and $\xi^{h}_{2}$ denote the first and the second hands’ joint states (e.g., the palm and fingers), and $o_{1}$ and $o_{2}$ represents the $6D$ poses of either two objects (e.g., stacking two cups), or two parts of the same object (e.g., bottle and bottle cap). The action at Step $t$ is given by $a(t)=[\hat{\xi}^{b}_{1}(t),\hat{\xi}^{b}_{2}(t),\hat{\xi}^{h}_{1}(t),\hat{\xi}^{h}_{2}(t)]$, composed of the target base poses ($\hat{\xi}^{b}_{1}$ and $\hat{\xi}^{b}_{2}$) and target joint positions ($\hat{\xi}^{h}_{1}$ and $\hat{\xi}^{h}_{2}$) for both hands. These actions are fed to a PD tracking controller to actuate both hands.
Note that our primary contributions pertain to the interaction phase of bimanual dexterous manipulation, in which two dexterous multi-fingered hands coordinate to complete the task after having grasped the necessary object(s). Most existing works focus solely on the interaction phase [40, 5, 41]. In Section. 3.6, we discuss how our approach can be extended to also tackle the acquisition phase, in which the hands learn to grasp objects before coordinating.
3.2 A Monolithic Approach
We begin by discussing the most straightforward approach one could take: a monolithic policy that utilizes all accessible environment states to plan actions for both hands: $\pi_{\mathrm{naive}}(\hat{\xi}^{b}_{1}(t),\hat{\xi}^{b}_{2}(t),\hat{\xi}^{h}_{1}(t),\hat{\xi}^{h}_{2}(t)|\xi^{b}_{1}(t),\xi^{b}_{2}(t),\xi^{h}_{1}(t),\xi^{h}_{2}(t),o_{1}(t),o_{2}(t))$. Training such a policy can be highly inefficient due to the high dimensionality of the observation and action spaces. Importantly, this monolithic implementation does not exploit the natural asymmetry found in most bimanual tasks. Below, we explain how AsymDex incorporates this insight.
3.3 Incorporating Asymmetry
When humans execute bimanual manipulation tasks, we tend to use our dominant hand to perform precise manipulations, and use a non-dominant hand to facilitate such manipulation [16, 17, 18]. For instance, when opening a bottle, we typically use our facilitating hand to move and reorient the bottle such that the bottle cap is closer to and oriented toward the dominant hand, which will then grasp the cap and uncap the bottle. During this cooperation process, our non-dominant hand moves and rotates the object with a firm grasp to facilitate the object manipulation by the dominant hand. Motivated by this, we assign different roles to each robot hand (i.e., a facilitating hand and a dominant hand) during bimanual manipulation.
Further, we make the observation that there tends to be no relative motion between the facilitating hand and the grasped object since the facilitating hand need only hold, move, and reorient the object (i.e., no in-hand reorientation). On the contrary, the hand dominant can interact freely with the object either directly or via another object. This observation suggests that the asymmetric manipulation and coordination strategy is neither dependent on nor influences the hand joints of the facilitating hand. As such, we can considerably reduce the observation and action spaces by accounting for asymmetry. Precisely, this observation allows us to define an asymmetric bimanual policy: $\pi_{\mathrm{asym}}(\hat{\xi}^{b}_{d}(t),\hat{\xi}^{b}_{f}(t),\hat{\xi}^{h}_{d}(t)|\xi^{b}_{d}(t),\xi^{b}_{f}(t),\xi^{h}_{d}(t),o_{1}(t),o_{2}(t))$ which use the facilitating hand to only to reposition and reorient the object. This considerable reduction in the dimensions of the observation and action spaces is likely to result in improved sample efficiency.
3.4 Incorporating Relative Observation and Action Spaces
1 Randomly initialize the two hand bases’ poses $\xi^{b}_{f}(0)$ and $\xi^{b}_{d}(0)$, and initialize object poses $o_{f}(0)$ and $o_{d}(0)$ based on $\xi^{b}_{f}(0)$ and $\xi^{b}_{d}(0)$. Initialize policy $\pi_{\theta}$.
2 for $iter\in\{1,...,\max\}$ do
3       Initialize replay buffer $\mathcal{B}=\varnothing$ ;
4       for $t\in\{1,...,M\}$ do
5             Simulate:
6             Collect hand and object states $\xi^{b}_{f}(t)$, $\xi^{b}_{d}(t)$, $\xi^{h}_{d}(t)$, $o_{f}(t)$, $o_{d}(t)$;
7             Compute relative states $\xi^{b}_{r}(t)=\xi^{b}_{a}(t){\circleddash}P_{f}$, $o_{r}(t)=o_{d}(t){\circleddash}P_{f}$;
8             Policy $\pi_{\mathrm{AsymDex}}(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t)|\xi^{b}_{r}(t),\xi^{h}_{d}(t),o_{r}(t))$ outputs relative actions;
9             Bimanual controller (Eqn. 1) computes $\hat{\xi}^{b}_{f}(t)$ and $\hat{\xi}^{b}_{d}(t)$ based on $\hat{\xi}^{b}_{r}(t)$;
10             if Meet reset condition then
11                  Reset environment;
12             end if
13            Environment physics steps with $\hat{\xi}^{b}_{f}(t),\hat{\xi}^{b}_{d}(t),\hat{\xi}^{h}_{d}(t)$;
14             Evaluate:
15             Compute reward $r(t)$
16             Collect observations $(\xi^{b}_{r}(t),\xi^{h}_{r}(t),o_{r}(t))$, actions $(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{r}(t))$, and reward $r(t)$ into buffer $\mathcal{B}$;
17
18       end for
19      Update the Policy $\pi_{\mathrm{AsymDex}}$ based on $\mathcal{B}$;
20
21 end for
Return: Trained policy $\pi_{\mathrm{AsymDex}}$
Algorithm 1 AsymDex: Interaction Phase
In addition to asymmetry, a key characteristic of bimanual dexterous manipulation is the synchronized motion of the two hands in which each hand moves in response to the other. In this section, we explain how we can further reduce the size of the observation and action spaces by defining relative and object-centric spaces that capture the relationships between the motions of two hands and the object(s) being manipulated. Indeed, the use of relative state spaces has shown to considerably benefit bimanual manipulation with simple end effectors [42, 43, 44, 23]. Furthermore, some of these prior work is limiting the relative space in a one-degree-of-freedom (1-DoF) action space [23], while AsymDex allows for complete 6-DoFs relative space.
Let $o_{f}=o_{1}$ be the state of the object being held by the facilitating hand, and let $o_{d}=o_{2}$ be the state of the object being manipulated by the dominant hand. We attach a coordinate frame to the object being held by the facilitating hand: $P_{f}$. Now, we can transform the observations $(\xi^{b}_{d}(t),\xi^{b}_{f}(t),\xi^{h}_{d}(t),o_{f}(t),o_{d}(t))$, which where originally defined in the world coordinate frame $P_{W}$, into the new coordinate frame $P_{f}$. Note that since there is no relative motion between the facilitating hand and the object that it’s holding, neither $o_{f}(t)$ nor $\xi^{b}_{f}(t)$ do not change in $P_{f}$. As such, both $o_{f}(t)$ and $\xi^{b}_{f}(t)$ can be neglected without losing any information. Now, transforming the remaining observations $(\xi^{b}_{d}(t),\xi^{h}_{d}(t),o_{d}(t))$ into the new coordinate $P_{f}$ yields $(\xi^{b}_{r}(t),\xi^{h}_{d}(t),o_{r}(t))$, respectively. Here, $\xi^{b}_{r}(t)$ denotes the 6D pose of the dominant hand base and $o_{r}(t)$ denotes the 6D pose of the object being manipulated by the dominant hand, both now defined relative to the object being held by the facilitating hand. Note that since $\xi^{h}_{d}(t)$ denotes the dominant hands’ joint states, it is not impacted by the change of coordinates. Similarly, we apply the same modifications to the action spaces of the asymmetry bimanual policy. This allows us to reduce the actions from $(\hat{\xi}^{b}_{d}(t),\hat{\xi}^{b}_{f}(t),\hat{\xi}^{h}_{d}(t))$ to $(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t))$, where $\hat{\xi}^{b}_{r}(t)$ is the target relative pose of the dominant hand now defined relative to the object being held by the facilitating hand.
Incorporating the above change of coordinates in addition to leveraging asymmetry, allows us to define AsymDex’s policy as $\pi_{\mathrm{AsymDex}}(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t)|\xi^{b}_{r}(t),\xi^{h}_{d}(t),o_{r}(t))$. Note that our formulation has significantly reduced the dimensions of both the state and action spaces, compared to the naive policy $\pi_{\mathrm{naive}}$ as defined in Section 3.2. See Appendix. A for RL algorithm and policy architecture.
3.5 Relative Pose Controller
To control the hand bases based on the target relative pose $\hat{\xi}^{b}_{r}(t)$ provided by $\pi_{\mathrm{AsymDex}}$, we designed a bimanual controller that computes both the target dominant hand base pose $\hat{\xi}^{b}_{d}(t)$ and the target facilitating hand base pose $\hat{\xi}^{b}_{f}(t)$ as follows
$\begin{split}\hat{\xi}^{b}_{d}(t)&=\alpha R^{o_{f}}_{world}\cdot dist(\hat{\xi}^{b}_{r}(t),\xi^{b}_{r}(t))+\xi^{b}_{d}(t),\\ \hat{\xi}^{b}_{f}(t)&=(\alpha-1)R^{o_{f}}_{world}\cdot dist(\hat{\xi}^{b}_{r}(t),\xi^{b}_{r}(t))+\xi^{b}_{f}(t),\end{split}$ (1)
where $R^{o_{f}}_{world}$ denotes the rotational transformation from Frame $P_{f}$ to the world frame $P_{W}$, $dist(\cdot)$ denotes the difference between two 6D poses. and $\alpha$ is a hyperparameter that controls the involvement of each hand. Throughout our experiments, we used fixed $\alpha=0.5$. The pseudo-code of the training process is included in Alg. 1.
3.6 Acquiring Objects
Figure 2: We decompose asymmetric bimanual dexterous manipulation into two phases: An acquisition phase and an interaction phase. We show that AsymDex can be integrated with learned grasping policies in order to seamlessly execute both phases.
While our approach as explained thus far deals with the challenge of coordinating two hands to accomplish asymmetric dexterous manipulation tasks, it assumes the task begins with that the object(s) of interest have already been grasped. However, in practice, robots must be able to tackle the challenge of grasping the necessary objects before the interaction between the two hands and the objects can begin. We refer to this phase as the acquisition phase. Most recent works on bimanual dexterous manipulation often entirely ignore the acquisition phase and focus purely on the interaction phase [7, 5, 40]. In contrast, we demonstrate that our approach can seamlessly accommodate the acquisition phase by i) leveraging the observation that the acquisition phase doesn’t require the coordination of two arms, and ii) employing recent advances in learning to grasp. Specifically, we demonstrate that we can seamlessly integrate AsymDex with PDGM [45], which can efficiently learn multi-fingered grasping policies by leveraging pre-grasp poses (see Fig. 2). Details about the grasping reward design are available in Appendix. B. We begin by executing the grasping policy in isolation and then ”turn on” the asymmetric policy learned by AsymDex after the object has been firmly grasped by the facilitating hand. If the task requires the dominant hand to also grasp a second object, we employ the same method to train a grasping policy for dominant hand to acquire the object, but switch the control of the dominant hand’s joints over the asymmetric policy after the object has been grasped.
4 Evaluation
We evaluated AsymDex on four asymmetric bimanual manipulation tasks and compared its performance against strong baselines that challenge our key design choices. We begin by explaining the various aspects of our experimental setup and follow with a discussion of results.
Platform: We conducted all our experiments using two ShadowHands – each a 30-DoF simulated multi-finger hand system (24-DoF hand + 6-DoF floating wrist base) built with Isaac Gym [46].
Tasks: We evaluated AsymDex and the baselines on the following four biannual manipulation tasks which contain both original (Block in cup, Bottle cap) and adapted tasks (Stack, Switch) from BiDexHand [20] (see Fig. 3 for visualization).
  • $\bullet$
    Block in cup: The two hands must coordinate to ensure that one hand places a block inside a cup that is being held by the other without letting either the cup or the block fall to the ground.
  • $\bullet$
    Stack: Two cups need to be stacked together. Each hand must hold a cup, and both must coordinate such that the two cups are aligned as one slides into the other.
  • $\bullet$
    Bottle cap: One hand must hold and reorient a bottle such that the other hand can grasp and separate the bottle cap from the bottle.
  • $\bullet$
    Switch: One hand holds and reorients a switch in a way that allows the other hand to turn it on.
Note that the hands have to coordinate and perform all four tasks without relying on a support surface such as a table and ensure that the object(s) are not dropped. See Appendix. C for details on state space design, sampling procedure, success criteria, and reward design for each task.
Figure 3: We evaluate AsymDex on four bimanual dexterous manipulation tasks.
Metrics: We quantify performance in terms of i) Task success rate (see Appendix. C for criteria) and ii) sample efficiency. We report both metrics across five random seeds in all experiments.
We evaluated AsymDex with two sets of experiments as described below.
4.1 Learning Bimanual Coordination
In this experiment, we focus on AsymDex’s effectiveness during the interaction phase. We initialize the environment such that the hands are at a pre-grasp pose around the objects using appropriate initial grasps (After the first timestep of the environment, the hand need to learn to catch and grasp the object from pre-grasp pose if it is not a facilitating hand). Note that this is a common assumption in recent methods that learn bimanual dexterous manipulation skills [5, 40]. Further, this allows us to isolate and examine AsymDex’s ability to learn to coordinate two multi-fingered hands. See Section 4.2 for the second experiment in which we also consider the challenge of acquiring the objects from a tabletop surface before interaction begins.
We compare AsymDex policy against the following baselines:
  • $\bullet$
    Monolithic: This policy doesn’t make assumptions about the structure of bimanual manipulation (see Sec. 3.2). As such, this baseline allows us to examine the necessity and effectiveness of leveraging both the asymmetry in hand roles and the relative action and observation spaces.
  • $\bullet$
    Asym-w/o-rel: This policy leverages asymmetry in hand roles, but learns over absolute observation and action places (see Sec. 3.3). As such, this baseline allows us to examine the necessity and effectiveness of relative action and observation spaces.
  • $\bullet$
    Rel-w/o-asym: This policy leverages the relative observation and action places, but ignores asymmetry. As such, it allows us to examine the necessity and effectiveness of asymmetry.
Figure 4: AsymDex consistently outperforms the baselines in terms of sample efficiency and success rate. Solid lines indicate mean trends and shaded areas show $\pm$ std. dev., over five random seeds.
We report the learning curves in Fig. 4 and success rates in Table 1. As can be readily observed, AsymDex is the only method that consistently performs well across all tasks, either performing comparably or outperforming all the baselines. While Rel-w/o-asym performs comparably to AsymDex on Stack, it is not able to match AsymDex’s performance on the other tasks. Both AsymDex and Rel-w/o-asym perform substantially better than the other two baselines across all tasks except Switch, where AsymDex performs much better than all three baselines. In fact, Asym-w/o-rel is able to learn an effective policy in only one task (Bottle cap), while Monolithic struggles on all four tasks. Curiously, Monolithic outperforms asym-w/o-rel on the Stack task. Qualitative analysis of rollouts reveals that, unlike Asym-w/o-rel, the Monolithic policy learns to use the facilitating hand’s fingers to orient the cup towards the bottom of the other cup.
Table 1: Success rates (mean $\pm$ std. dev.) in Interaction phase
Task Method Monolithic Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
Block in cup $0.0429\pm 0.0266$ $0.0164\pm 0.0190$ $0.3052\pm 0.1577$ 0.7701 $\pm$ 0.0559
Stack $0.2517\pm 0.2809$ $0.0044\pm 0.0037$ 0.9443 $\pm$ 0.0136 $0.9233\pm 0.0250$
Bottle cap $0.6125\pm 0.1734$ $0.7893\pm 0.1277$ $0.8073\pm 0.1389$ 0.8301 $\pm$ 0.1797
Switch $0.0563\pm 0.0126$ $0.1626\pm 0.0882$ $0.1149\pm 0.0176$ 0.6700 $\pm$ 0.0359
Taken together, the above observations reveal a few insights. First, when used in isolation, neither asymmetry nor relative motion are sufficient across all tasks. In particular, while each might provide sufficient structure to accomplish some tasks, they prove to be less effective on other tasks. Second, the use of relative motion offers a larger boost in performance compared to asymmetry, likely due to the fact that relative spaces avoid unnecessary exploration (e.g., when the two hands move in parallel) while allowing the facilitating hand to exhibit more complex behaviors. Third, ignoring both asymmetry and relative motion hardly leads to success.
4.2 Learning to Grasp and Coordinate
In this experiment, we evaluate AsymDex’s ability to incorporate the object acquisition phase in addition to the interaction phase. Specifically, we initialize the environment for each task such that the objects of interest are placed on a tabletop surface. As such, each method needs to learn both to grasp the necessary objects and to coordinate the two hands to complete the tasks.
For AsymDex, we follow the same strategy as introduced in Section 3.6. We compare AsymDex’s performance against the following baselines:
  • $\bullet$
    1-stage-monolithic: This baseline uses a single policy to learn both the grasping and interaction phases for both hands, and thus allows us to investigate the benefits of AsymDex’s two-phase decomposition.
  • $\bullet$
    2-stage-monolithic: This policy benefits from the same phase decomposition as AsymDex policy, but leverages neither asymmetry nor relative motion. As such, this baseline allows us to examine if this task can be solved merely with two-phase decomposition.
To ensure a fair comparison, we provide pre-grasp pose annotations to both baselines. Further, we ensure that the total number of env. interactions (the number of one stage or the sum of two stages) is the same across AsymDex and the baselines. See Appendix. B for details of the grasping learning.
We report the overall roll-out success rate of all methods for two tasks across five random seeds in Table. 2. We find that AsymDex significantly outperforms the other two baselines in both tasks, suggesting that combining phase decomposition with AsymDex’s other two design choices (asymmetry and relative state) results in policies that can effectively handle both the acquisition and the interaction phases of bimanual dexterous manipulation. The fact that 2-stage-monolithic baseline outperforms the 1-stage-monolithic baseline demonstrates the inherent benefits of phase decomposition. Our qualitative analysis of Block in cup task revealed that 1-stage-monolithic policy learns to tip the cup over and push the block towards the cup. In contrast, both the two-stage policies learn the expected behavior. This suggests that the phase decomposition nudges the grasping and interaction policies to learn reasonable behaviors that complement each other.
Table 2: Success rates (mean $\pm$ std. dev.) after combining acquisition and interaction phases
Task Method 1-stage-monolithic 2-stage-monolithic 2-stage-AsymDex
Block in cup $0.0321\pm 0.0251$ $0.1505\pm 0.1059$ 0.7938 $\pm$ 0.0897
Bottle cap $0.1680\pm 0.2695$ $0.6407\pm 0.1141$ 0.8726 $\pm$ 0.0600
5 Conclusion
Our framework (AsymDex) is capable of learning complex asymmetric bimanual dexterous manipulation tasks via reinforcement without relying on demonstrations. We introduced and validated the need for AsymDex’s two crucial ingredients: assigning asymmetric roles to the two hands, and using relative observation and action spaces. Our evaluation results reveal that the combination of these choices consistently leads to better sample efficiency and success rates across different tasks.
6 Limitations and Future Works
Our work has revealed a number of limitations and avenues for future research. First, AsymDex in its current form cannot handle certain bimanual tasks that require complex multi-finger manipulation from both hands (e.g., reorienting a heavy object, dynamic handover). Second, AsymDex does not consider the kinodynamic constraints that might result from manipulator arms. Third, AsymDex has not yet been evaluated on hardware. Fourth, behaviors produced by AsymDex are not always natural or human-like due to lack of necessary incentives.
Appendices
Appendix A RL Training
We use Proximal Policy Optimization (PPO) [47] algorithm to train all policies $\pi_{\mathrm{naive}},\pi_{\mathrm{asym}},$ and $\pi_{\mathrm{AsymDex}}$ with their corresponding value functions. Both policies and value functions are parameterized via a three-layer MLP network. The size of hidden layers for each is i) policy: (256, 256, 128), ii) value function: (512, 512, 512). The activation functions are all set as Exponential Linear Unit (ELU). We use the same PPO hyperparameters for all the baselines and AsymDex ($learning\ rate:3\times 10^{-4},\gamma:0.98,\lambda:0.95$). We train the polices on a computer with a single Nividia RTX 4090 GPU.
Appendix B Grasping Learning
Two-stage policy
Both AsymDex (our approach) and 2-stage-monolithic policy (one of the baselines in Sec. 4.2) are two-stage policies. Therefore, they can first learn a grasping policy for the facilitating hand (or two grasping policies for facilitating hand and dominant hand respectively). Such policy $\pi_{\mathrm{grasp}}(\hat{\xi}^{h}_{f}(t)|\xi^{h}_{f}(t),\xi^{b}_{f}(t)\circleddash X_{o_{f}})$ takes in the hand joint states and the relative pose between hand and the object, and outputs the target hand joint positions to grasp the object firmly. We first provide pre-grasp annotations [45], which allows the hands to initialize at the position close to the objects with proper joint positions. Then we script the $6D$ lifting hand base motions and design the the following rewards, which is the same across all objects.
$Reward=R_{rel\ pos}+R_{rel\ rot}$
The relative position reward $R_{rel\ pos}=(\alpha-||x_{obj}-x_{initial}||)*\beta$, where $x_{obj}$ is the current relative position between the object and the hand, and $x_{initial}$ is the initial relative position between the object and the hand. The $\alpha,\beta\in\mathcal{R}_{+}$ are hyper-parameters. The relative rotation reward $R_{rel\ pos}=<u_{obj},u_{hand}>$, where $u_{obj}$ is the object direction vector, $u_{hand}$ is the hand direction vector, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the inner product of two vectors. We define the object direction vector and hand direction vector to be the same at the beginning of the grasping phase. Both rewards encourage the hand to keep a constant relative pose, i.e., grasping the object, during the script motion.
One-stage policy
Another baseline in Sec. 4.2, i.e., the 1-stage-monolithic policy, does not incorporate the task decomposition. Therefore, it only uses the task-specific interaction rewards (see Appendix. C) to learn how to complete the entire bimanual task. For a fair comparison, both hands also start at the pre-grasp poses.
Appendix C Task Design
In this section, we show the details for each task.
State Space Design
For each task, the hand joint states $\xi^{h}_{f}(t)$, $\xi^{h}_{d}(t)$ include the 24-DoF hand joint positions and the 24-DoF hand joint velocities. We use quaternions to represent rotation part of object and hand base poses. And for all policies, we also include the previous actions in the policy input. For the block in cup task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the cup and the block respectively. For the stack task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of two cups. For the Bottle cap task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the bottle and the cap respectively. For the Switch task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the switch body and the button respectively. The dimensions of the observation and action spaces of each policy are shown in Table. 3. It is obvious that AsymDex policy significantly reduces the state dimensions.
Table 3: Dimension of Observation and action spaces. For all tasks, the dimensions are identical.
Monolithic Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
Observation $176$ $108$ $163$ $88$
Action $52$ $32$ $46$ $26$
Sampling Procedure
  • Place block in cup: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.57,1.57)$, in radians. The block is initialized in the dominant hand. Thus its position and rotation is calculated based on the initial position and rotation of dominant hand base. The initial position of the facilitating hand base is at $[0.55,0.6,0.8]$.
  • Stack cups: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.57,1.57)$, in radians. The cup is initialized in the dominant hand. Thus its position and rotation is calculated based on the initial position and rotation of dominant hand. The initial position of the facilitating hand base is at $[0.55,0.6,0.8]$.
  • Open bottle cap: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.58,0.62)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.21,-0.19)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.58,0.62)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.0,1.0)$, in radians. The initial position of the facilitating hand base is randomized: $x_{f}\in\mathcal{X}\sim\mathcal{U}(0.53,0.57)$, $y_{f}\in\mathcal{Y}\sim\mathcal{U}(0.59,0.61)$, $z_{f}\in\mathcal{Z}\sim\mathcal{U}(0.43,0.45)$. For the rotation of the facilitating hand, we randomly rotate it around the axis along the arm at a random angle, $\beta\in\mathcal{B}\sim\mathcal{U}(-0.5,0.5)$, in radians. The bottle is initialized in the facilitating hand. Thus its position and rotation is calculated based on the initial position and rotation of facilitating hand base.
  • Turn on switch: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.2,0.6)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.25,-0.05)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.5,0.9)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.0,1.0)$, in radians. The initial position of the facilitating hand base is randomized: $x_{f}\in\mathcal{X}\sim\mathcal{U}(0.2,0.6)$, $y_{f}\in\mathcal{Y}\sim\mathcal{U}(0.05,0.25)$, $z_{f}\in\mathcal{Z}\sim\mathcal{U}(0.41,0.81)$. For the rotation of the facilitating hand, we randomly rotate it around the axis along the arm at a random angle, $\beta\in\mathcal{B}\sim\mathcal{U}(-1.0,1.0)$, in radians. The switch is initialized in the facilitating hand. Thus its position and rotation is calculated based on the initial position and rotation of facilitating hand base.
Success Criteria
  • Place block in cup The task is considered successful if the distance of the block center and the cup center is smaller than 0.035 meters. This distance makes sure the task is only considered successful when the block is inside the cup. If the block falls on the ground or has not entered the cup within a certain time step, the task is considered failed.
  • Stack cups The task is considered successful if the distance between the cup centers is smaller than 0.02 meters. If either cup falls on the ground or has not been stacked within a certain time step, the task is considered failed.
  • Open bottle cap The task is considered successful if the cap is taken off from its original position 0.05 meters away within a time duration, and is considered failed otherwise.
  • Turn on switch The button and the switch body are connected by a revolute joint ranging from 0 to 0.5585 rads. The task is considered successful if the button is pressed and rotated 0.3585 rads within a time duration, and is considered failed otherwise.
Reward Design
The reward design is similar across all tasks:
$Reward=R_{hand\ distance}+R_{progress}+R_{action\ penalty}+R_{success\ bonus}$
For each task, $R_{action\ penalty}=-||a(t)||^{2}$, and the $R_{success\ bonus}$ is the task success reward. $R_{hand\ distance}$ and $R_{progress}$ are slightly different for each tasks.
  • Place block in cup: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth. $R_{progress}=-||x_{cup}-x_{block}||$, where $x_{cup}$ is the position of the cup, and $x_{block}$ is the position of the block.
  • Stack cups: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth, which is grasped by the facilitating hand. $R_{progress}=-||x_{cupd}-x_{cupf}||$, where $x_{cupf}$ is the position of the cup grasped by the facilitating hand, and $x_{cupd}$ is the position of the cup grasped by the dominant hand.
  • Open bottle cap: $R_{hand\ distance}=(1-(||x_{index}-x_{cap}||+||x_{thumb}-x_{cap}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{cap}$ is the position of the bottle cap. $R_{progress}=||x_{cap}-x_{bottle\ top}||$, where $x_{cap}$ is the position of the cap, and $x_{bottle\ top}$ is the position of the top of the bottle.
  • Turn on switch: $R_{hand\ distance}=(1-(||x_{index}-x_{button}||+||x_{thumb}-x_{button}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{button}$ is the position of the button. $R_{progress}=2*\theta_{button}$, where $\theta_{button}$ is the rotated angle of the joint that connects the button and the switch body.

### v2

@@
-AsymDex: Leveraging Asymmetry and Relative Motion in Learning Bimanual Dexterity
+AsymDex: Asymmetry and Relative Coordinates for RL-based Bimanual Dexterity
 Abstract
-We present Asymmetric Dexterity (AsymDex), a novel reinforcement learning (RL) framework that can efficiently learn asymmetric bimanual skills for multi-fingered hands without relying on demonstrations, which can be cumbersome to collect.
-Two crucial ingredients enable AsymDex to reduce the observation and action space dimensions and improve sample efficiency.
-First, AsymDex leverages the natural asymmetry found in human bimanual manipulation and assigns specific and interdependent roles to each hand: a facilitating hand that moves and reorients the object, and a dominant hand that performs complex manipulations on said object.
-Second, AsymDex defines and operates over relative observation and action spaces, facilitating responsive coordination between the two hands.
-Further, AsymDex can be easily integrated with recent advances in grasp learning to handle both the object acquisition phase and the interaction phase of bimanual dexterity.
-Unlike existing RL-based methods for bimanual dexterity, which are tailored to a specific task, AsymDex can be used to learn a wide variety of bimanual tasks that exhibit asymmetry.
-Detailed experiments on four simulated asymmetric bimanual dexterous manipulation tasks reveal that AsymDex consistently outperforms strong baselines that challenge its design choices, in terms of success rate and sample efficiency.
-The project website is at https://sites.google.com/view/asymdex-2024/.
-Keywords: Dexterous Manipulation, Bimanual Manipulation
+We present Asymmetric Dexterity (AsymDex), a novel and simple reinforcement learning (RL) framework that can efficiently learn a large class of bimanual skills in multi-fingered hands without relying on demonstrations.
+Two crucial insights enable AsymDex to reduce the observation and action space dimensions and improve sample efficiency.
+First, true ambidexterity is rare in humans and most of us exhibit strong “handedness”.
+Inspired by this observation, we assign complementary roles to each hand: the facilitating hand repositions and reorients one object, while the dominant hand performs complex manipulations to achieve the desired result (e.g., opening a bottle cap, or pouring liquids).
+Second, controlling the relative motion between the hands is crucial for coordination and synchronization of the two hands.
+As such, we design relative observation and action spaces and leverage a relative-pose tracking controller.
+Further, we propose a two-phase decomposition in which AsymDex can be readily integrated with recent advances in grasp learning to facilitate both the acquisition and manipulation of objects using two hands.
+Unlike existing RL-based methods for bimanual dexterity with multi-fingered hands, which are either sample inefficient or tailored to a specific task, AsymDex can efficiently learn a wide variety of bimanual skills that exhibit asymmetry.
+Detailed experiments on seven asymmetric bimanual dexterous manipulation tasks (four simulated and three real-world) reveal that AsymDex consistently outperforms strong baselines that challenge our design choices.
+The project website is at https://sites.google.com/view/asymdex-2025/.
+Keywords: Dexterous Bimanual Manipulation, Multi-Fingered Hands
 1 Introduction
-We tackle the challenge of learning bimanual dexterous manipulation skills on multi-fingered hands using reinforcement learning (RL).
-Bimanual skills are crucial for robots operating in human environments as they allow for more complex and flexible manipulation compared to a single hand [1, 2, 3, 4, 5, 6, 7].
-While learning dexterous manipulation skills on a single hand presents numerous challenges [8, 9, 10, 11, 12, 13, 14], learning bimanual dexterity can be significantly more challenging due to the higher-dimensional state and action spaces and the need to coordinate and synchronize the movement of two hands [1].
-To circumvent the challenges introduced by the high-dimensional state and action spaces of bimanual dexterous manipulation, we take inspiration from how humans approach this challenge.
-Humans exhibit a natural asymmetry in how we use each of our hands when we perform most bimanual tasks.
-Specifically, we tend to use one hand to reposition and reorient an object being manipulated so as to make it easier for the other hand to achieve the desired manipulation objectives.
-While the asymmetric assumption might seem restrictive at first glance, rich bodies of work in human biomechanics and evolution reveal its significance and necessity [15, 16, 17, 18].
-Evolutionary biologists posit that such handedness evolved in humans and great apes to meet the escalating cognitive demands of tool use and complex manipulation [19].
-Indeed, a large class of real-world bimanual tasks admit this asymmetry (e.g., attachment, detachment, assembly, pouring).
-A key insight that we leverage is that the natural asymmetry between the two hands can reduce the dimensionality of the bimanual dexterous manipulation, and in turn improve effectiveness and sample efficiency.
-Using this insight, we contribute a novel learning framework, dubbed Asymmetric Dexterity (AsymDex), that can efficiently learn asymmetric bimanual dexterous manipulation tasks based on RL (see Fig.
-1 for a block diagram).
-Note that AysmDex does not require the cumbersome collection of demonstrations and can learn bimanual skills only using reinforcement learning.
-Figure 1: Our approach (AsymDex) efficiently learns asymmetric bimanual dexterous manipulation skills based on reinforcement learning by effectively leveraging i) the natural asymmetry in the hands’ roles and ii) relative state and action spaces that encourage synchronization.
-AsymDex has two crucial ingredients.
-First, it introduces asymmetry by defining a dominant hand and a facilitating hand.
-While the facilitating hand learns to reposition and reorient the object, the dominant hand learns complex manipulation skills (including in-hand manipulation).
-Note that there’s no relative movement between the grasped object and the facilitating hand.
-As a result, AsymDex holds the fingers of the facilitating hands in the grasping pose and only controls the 6D motion of its base.
-On the other hand, AsymDex learns to control both the base and fingers of the dominant hand.
-Second, AsymDex reasons about and controls the relative motion between the dominant and facilitating hands.
-We define relative observation and action spaces that incentivize responsive coordination between the two hands without resorting to explicit time-dependence.
-We also leverage the observation that bimanual manipulation in practice is composed of two distinct phases: i) the acquisition phase in which objects are grasped from surfaces, and ii) the interaction phase where the two hands coordinate to perform the bimanual task.
-Unlike many existing methods that ignore the acquisition phase, we show that this decomposition enables AysmDex to be seamlessly integrated with learned grasping policies to enable fluent execution.
-In summary, we contribute AsymDex – a novel framework for asymmetric bimanual dexterity that leverages the natural asymmetry in hand roles and relative observation and action spaces.
-We evaluate AsymDex on four complex bimanual dexterous tasks (adapted from BiDexHand [20]) and compare against strong baselines that challenge its design choices.
-Our results show that AsymDex consistently outperforms the baselines in terms of success rate and sample efficiency.
+Bimanual dexterity is crucial for robots operating in human environments as they allow for complex yet flexible manipulation compared to a single hand [1, 2, 3, 4, 5, 6].
+We are interested in learning a wide range of bimanual dexterous skills on multi-fingered hands purely from reinforcement.
+While learning on a single multi-fingered hand is known to be challenging [7, 8, 9, 10, 11, 12, 13], learning bimanual dexterity is made more challenging due to the higher-dimensionality and the need for coordination and synchronization of two hands [1, 14].
+These challenges are only exacerbated when these skills have to be learned from reinforcement, explaining why existing efforts often either resort to expert demonstrations [15, 16, 17, 18] or limit themselves to specific tasks [19, 20].
+We rely on two insights to tackle the challenges of bimanual dexterity.
+First, we are inspired by how humans and other great apes approach these challenges: there is a natural asymmetry in how we use each of our hands when we perform most bimanual tasks [21].
+Specifically, we tend to use one hand to reposition and reorient an object so as to make it easier for the other hand to perform complex manipulation.
+While leveraging such asymmetry might appear to restrict the class of bimanual skills we can learn, rich bodies of work in human biomechanics and evolution reveal its prevalence and necessity [21, 22, 23, 24].
+Evolutionary biologists posit handedness evolved to meet the escalating cognitive demands of tool use, language, and complex manipulation [25].
+Indeed, a large number of real-world tasks admit such asymmetry (e.g., attachment, detachment, assembly, and pouring).
+Looking closely at the asymmetry in bimanual dexterity reveals our second insight.
+Our non-dominant hand tends to hold an object firmly as we reorient and reposition it relative to the dominant hand or the object being held by the dominant hand (e.g., tilting a pen before uncapping).
+This suggests that there is often little to no in-hand movement of the object grasped by the non-dominant hand, and robust synchronization can be achieved by ensuring relative movement of the two hands.
+Figure 1: Our approach (AsymDex) efficiently learns asymmetric bimanual dexterous manipulation skills based on reinforcement learning by effectively leveraging i) the natural asymmetry in the hands’ roles and ii) relative state and action spaces that prioritize synchronization.
+We contribute a novel RL-based learning framework for bimanual dexterity, dubbed Asymmetric Dexterity (AsymDex) by operationalizing the above two insights (see Fig.
+3).
+To incorporate asymmetry, we define a dominant hand and a facilitating hand.
+While the dominant hand learns complex skills across all its degrees of freedom, the facilitating hand learns to reposition and reorient the object by controlling the 6D pose of its base (i.e., no finger movement).
+This allows us to both reduce the dimensionality of the observation and action spaces and tightly integrate the roles of the two hands.
+To ensure coordination and synchronization, AsymDex operates over relative observation and action spaces that incentivize flexible coordination of the two hands without resorting to explicit time-dependence or task-specific coordinate frame designs.
+We also leverage the observation that bimanual manipulation in practice is composed of two distinct phases: i) the acquisition phase in which objects are grasped, and ii) the interaction phase in which the two hands coordinate to perform the bimanual task.
+Unlike many existing methods that entirely ignore the acquisition phase[2, 14, 19, 26], we show that this decomposition enables AysmDex to be seamlessly integrated with learned grasping policies to enable fluent execution.
+In summary, we contribute AsymDex – a novel Rl-based framework for learning a wide variety of bimanual dexterous skills by taking inspiration from two key aspects of human bimanual dexterity: i) asymmetric hand roles, and ii) relative hand movement.
+We conduct comprehensive experiments on seven complex tasks (four simulated and three real) and compare against strong baselines that challenge the need for AsymDex’s structural inductive biases.
+Our results show that AsymDex consistently outperforms these baselines in terms of both task performance and sample efficiency.
 2 Related Work
-In this section, we contextualize our contributions within relevant sub-fields.
 Learning Bimanual Manipulation: Several existing methods focus on learning bimanual skills, but are often limited to simple end-effectors.
@@
 Recent works have tackled these challenges using imitation learning (IL) or reinforcement learning (RL) and demonstrate impressive performance [7, 8, 9, 10, 11, 13, 39, 40, 41, 42].
-However, IL-based methods, including those that combine RL and IL [8, 10], rely either on complex infrastructure and retargeting methods to collect demonstrations  [8, 35, 14, 38, 39] or pre-trained expert policies [11, 12, 37].
-On the other hand, RL-based methods do not share these constraints as they learn skills via reinforcement.
-However, RL methods tend to require single amounts of exploration even for unimanual dexterous manipulation [9, 34, 10, 36].
+However, IL-based methods rely either on complex infrastructure and retargeting methods to collect demonstrations  [7, 13, 40, 43, 44] or pre-trained expert policies [10, 11, 42].
+On the other hand, RL-based methods do not share these constraints as they learn skills via reinforcement, but tend to require significant exploration even for dexterous manipulation with a single hand [8, 9, 39, 41].
 As we show in our experiments, naive application of RL-based methods is not effective for bimanual dexterous manipulation due to the increased dimensionality and the need for coordination.
-We also show that AsymDex is able to efficiently learn complex bimanual dexterous manipulation skills using RL without relying on demonstrations.
 Learning Bimanual Dexterous Manipulation: A few recent studies have focused on learning bimanual dexterity.
@@
 While these methods are specifically designed to solve a particular task, AsymDex is capable of efficiently learning different bimanual dexterous manipulation tasks.
-3 Learning Asymmetric Bimanual Dexterous Manipulation Skills
-In this section, we formulate the problem of asymmetric bimanual dexterous manipulation and introduce the different elements of our approach (AsymDex).
-3.1 Preliminaries
-Consider the problem of bimanual dexterous manipulation, in which two multi-fingered hands coordinate to manipulate up to two objects.
+Some recent studies investigate generalized learning method by utilizing expert demonstration for efficient RL training [15, 16, 17, 18], while AsymDex can efficiently learn bimanual manipulation skills without relying on demonstration.
+3 Problem Formulation
+We first formulate the general problem of bimanual dexterous manipulation, and then introduce asymmetric bimanual dexterity.
+Consider the general problem of bimanual dexterous manipulation, in which two multi-fingered hands coordinate to manipulate up to two objects.
 Formally, this problem can be defined as a Partially-Observable Markov Decision Process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{Z},\mathcal{A},\mathcal{R},\mathcal{P})$, where $\mathcal{S}\in\mathbb{R}^{n}$ is the state space, $\mathcal{Z}\in\mathbb{R}^{m}$ is the observation space, $\mathcal{A}\in\mathbb{R}^{u}$ is the action space, ${\mathcal{R}:\mathbb{R}^{m}\times\mathbb{R}^{u}\rightarrow\mathbb{R}}$ is the reward function, and ${\mathcal{P}:\mathbb{R}^{n}\times\mathbb{R}^{u}\rightarrow\mathbb{R}^{n}}$ is the environment dynamics.
@@
 Given this formulation, the problem boils down to learning a policy $\pi:\mathcal{Z}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $E_{\pi}[\Sigma_{t=0}^{T-1}\gamma^{t}\mathcal{R}(z(t),a(t))]$.
-Observation and Action Space: The observation space $\mathcal{Z}$ is composed of hand and object measurements.
-At step $t$, $z(t)$ contains $\xi^{b}_{1}(t)\in\mathbb{R}^{6}$, $\xi^{b}_{2}(t)\in\mathbb{R}^{6}$, $\xi^{h}_{1}(t)\in\mathbb{R}^{n_{1}}$, $\xi^{h}_{2}(t)\in\mathbb{R}^{n_{2}}$, $o_{1}(t)\in\mathbb{R}^{6}$, and $o_{2}(t)\in\mathbb{R}^{6}$, where $\xi^{b}_{1}$ and $\xi^{b}_{2}$ denote the first and the second hand bases’ $6D$ poses, $\xi^{h}_{1}$ and $\xi^{h}_{2}$ denote the first and the second hands’ joint states (e.g., the palm and fingers), and $o_{1}$ and $o_{2}$ represents the $6D$ poses of either two objects (e.g., stacking two cups), or two parts of the same object (e.g., bottle and bottle cap).
-The action at Step $t$ is given by $a(t)=[\hat{\xi}^{b}_{1}(t),\hat{\xi}^{b}_{2}(t),\hat{\xi}^{h}_{1}(t),\hat{\xi}^{h}_{2}(t)]$, composed of the target base poses ($\hat{\xi}^{b}_{1}$ and $\hat{\xi}^{b}_{2}$) and target joint positions ($\hat{\xi}^{h}_{1}$ and $\hat{\xi}^{h}_{2}$) for both hands.
-These actions are fed to a PD tracking controller to actuate both hands.
-Note that our primary contributions pertain to the interaction phase of bimanual dexterous manipulation, in which two dexterous multi-fingered hands coordinate to complete the task after having grasped the necessary object(s).
+Observation and Action Spaces: The observation space $\mathcal{Z}$ is composed of hand and object states.
+At time step $t$, $z(t)=[\xi_{1}(t),\xi_{2}(t),\xi^{obj}(t)]$, where $\xi_{1}(t)$ contains the first hand’s current full (fingers + wrist) configuration $\xi^{h}_{1}(t)\in\mathbb{R}^{n_{1}}$ and the 6 DOF pose of its base $\xi^{b}_{1}(t)\in\mathbb{SE}(3)$, $\xi_{2}(t)$ contains the corresponding elements for the second hand, and $\xi^{obj}(t)$ contains the 6 DOF poses of either two objects (e.g., stacking two cups) or parts of one object (e.g., bottle and lid).
+The joint action at time step $t$ is given by $a(t)=[\hat{\xi}_{1}(t),\hat{\xi}_{2}(t)]$, where $\hat{\xi}_{1}(t)$ denotes the target joint configuration of the first hand $\hat{\xi}^{h}_{1}(t)$ and the target 6 DOF pose of its base $\hat{\xi}^{b}_{1}(t)$, and $\hat{\xi}_{2}(t)$ denotes the corresponding targets for the second hand.
+With the above definitions, we can now define the problem of learning bimanual dexterity as one of learning a monolithic symmetric policy: $\pi_{\mathrm{sym}}(a(t)|z(t))$.
+Asymmetric Dexterity Problem: Asymmetric dexterity can be viewed as a broad subclass to the above general class of problems.
+Inspired by strategies employed by humans and other great apes, we are interested in tasks in which one multi-fingered hand performs complex and precise manipulations while the other plays a facilitating role by supporting and actively reorienting objects of interest (e.g., stacking, attachment, detachment, etc.).
+Note that our formulation does not restrict the movement of the second hand; it merely restricts the relative movement between the second hand and the object being held.
+As explained in Sec.
+1, a large number of bimanual tasks exhibit asymmetry, hinting at handedness in most humans.
+We are thus interested in learning an asymmetric policy $\pi_{\mathrm{AsymDex}}(\cdot)$ with carefully-defined observation $\mathcal{Z}_{\mathrm{AsymDex}}$ and action $\mathcal{A}_{\mathrm{AsymDex}}$ spaces in an effort to improve both effectiveness and sample efficiency.
+Note that our primary contributions and the asymmetric assumption pertain to the interaction phase of bimanual dexterity, in which the two hands actively coordinate to complete the task after having grasped the necessary object(s).
 Most existing works focus solely on the interaction phase [19, 20, 45].
-In Section.
-3.6, we discuss how our approach can be extended to also tackle the acquisition phase, in which the hands learn to grasp objects before coordinating.
-3.2 A Monolithic Approach
-We begin by discussing the most straightforward approach one could take: a monolithic policy that utilizes all accessible environment states to plan actions for both hands: $\pi_{\mathrm{naive}}(\hat{\xi}^{b}_{1}(t),\hat{\xi}^{b}_{2}(t),\hat{\xi}^{h}_{1}(t),\hat{\xi}^{h}_{2}(t)|\xi^{b}_{1}(t),\xi^{b}_{2}(t),\xi^{h}_{1}(t),\xi^{h}_{2}(t),o_{1}(t),o_{2}(t))$.
-Training such a policy can be highly inefficient due to the high dimensionality of the observation and action spaces.
-Importantly, this monolithic implementation does not exploit the natural asymmetry found in most bimanual tasks.
-Below, we explain how AsymDex incorporates this insight.
+In Sec.
+4.4, we discuss how our approach can be readily extended to also tackle the acquisition phase (learning to grasp the necessary objects before coordinating).
+4 AsymDex: Learning Asymmetric Dexterity
+While one could learn the symmetric policy $\pi_{sym}(\cdot)$ as defined in Sec.
+3, training such a policy can be inefficient or ineffective due to the high dimensionality of the observation and action spaces (see Sec.
+5 for empirical evidence).
+Importantly, a symmetric approach ignores the natural asymmetry found in most bimanual tasks.
+Below, we explain how AsymDex overcomes these challenges.
 4.1 Incorporating Asymmetry
-When humans execute bimanual manipulation tasks, we tend to use our dominant hand to perform precise manipulations, and use a non-dominant hand to facilitate such manipulation [16, 17, 18].
-For instance, when opening a bottle, we typically use our facilitating hand to move and reorient the bottle such that the bottle cap is closer to and oriented toward the dominant hand, which will then grasp the cap and uncap the bottle.
-During this cooperation process, our non-dominant hand moves and rotates the object with a firm grasp to facilitate the object manipulation by the dominant hand.
-Motivated by this, we assign different roles to each robot hand (i.e., a facilitating hand and a dominant hand) during bimanual manipulation.
-Further, we make the observation that there tends to be no relative motion between the facilitating hand and the grasped object since the facilitating hand need only hold, move, and reorient the object (i.e., no in-hand reorientation).
-On the contrary, the hand dominant can interact freely with the object either directly or via another object.
-This observation suggests that the asymmetric manipulation and coordination strategy is neither dependent on nor influences the hand joints of the facilitating hand.
-As such, we can considerably reduce the observation and action spaces by accounting for asymmetry.
-Precisely, this observation allows us to define an asymmetric bimanual policy: $\pi_{\mathrm{asym}}(\hat{\xi}^{b}_{d}(t),\hat{\xi}^{b}_{f}(t),\hat{\xi}^{h}_{d}(t)|\xi^{b}_{d}(t),\xi^{b}_{f}(t),\xi^{h}_{d}(t),o_{1}(t),o_{2}(t))$ which use the facilitating hand to only to reposition and reorient the object.
-This considerable reduction in the dimensions of the observation and action spaces is likely to result in improved sample efficiency.
+Motivated by the natural asymmetry in human bimanual manipulation [22, 23, 24], we assign different roles to each robot hand during their interaction: a facilitating hand that is responsible for holding and repositioning and reorienting the object, and a dominant hand which is responsible for fine-grained dexterous manipulation of the object(s).
+We note that there tends to be no relative motion between the facilitating hand and the grasped object in asymmetric dexterity since the facilitating hand need only hold, move, and reorient the object (i.e., no in-hand reorientation).
+In contrast, the dominant hand can interact freely with the object(s).
+This suggests that the asymmetric dexterity is neither dependent on nor influences the finger joints of the facilitating hand during the interaction phase.
+As such, we can considerably reduce the observation and action spaces by defining an asymmetry-only bimanual policy: $\pi_{\mathrm{asym}}(a_{\mathrm{asym}}(t)|z_{\mathrm{asym}}(t))$ with actions $a_{\mathrm{asym}}(t)=[\hat{\xi}_{d}(t),\hat{\xi}^{b}_{f}(t)]$ and observations $z_{\mathrm{asym}}(t)=[\xi_{d}(t),\xi^{b}_{f}(t),\xi^{obj}(t)]$.
+This change both reduces the dimensionality and ensures that the facilitating hand doesn’t learn unproductive or unnecessary behaviors.
 4.2 Incorporating Relative Observation and Action Spaces
+Figure 2: AsymDex’s observation and Action Spaces.
+In addition to asymmetry, a key characteristic of bimanual dexterity is the synchronized and responsive movement of the two hands.
+We can further reduce the size of the observation and action spaces and introduce tight coupling between the hands’ behaviors by defining relative and object-centric coordinates that capture the relationships between the movements of two hands and the object(s) being manipulated.
+Indeed, the use of relative state spaces has shown to considerably benefit bimanual manipulation with simple end effectors [26, 46, 47, 48].
+While some of these prior works limit the relative space to a one-degree-of-freedom (1-DoF) action space [26], AsymDex allows for complete 6-DoFs relative space.
+Let $\xi^{obj}_{f}(t)$ be the state of the object being held by the facilitating hand, and let $\xi^{obj}_{d}(t)$ be the state of the object being manipulated by the dominant hand.
+We attach a coordinate frame to the object being held by the facilitating hand: $P_{f}$.
+Now, we can transform the asymmetry-only observations $z_{\mathrm{asym}}(t)$ (originally defined in the absolute or world coordinate frame) into the new coordinate frame $P_{f}$.
+Note that since there is no relative motion between the facilitating hand and the object that it is holding, neither $\xi^{obj}_{f}(t)$ nor $\xi^{b}_{f}(t)$ change in $P_{f}$, and can thus be removed from our observation space without losing any information.
+Now, transforming the observations ($\xi^{b}_{d}(t),\xi^{obj}_{d}(t)$) onto $P_{f}$ yields $z_{\mathrm{AsymDex}}=[\xi^{h}_{d}(t),\xi^{b}_{r}(t),\xi^{obj}_{r}(t)]$, where $\xi^{b}_{r}(t)$ and $\xi^{obj}_{r}(t)$ respective denote the 6D relative poses of the dominant hand base and the object being manipulated by the dominant hand, both defined with respect to the object being held by the facilitating hand.
+Note that since $\xi^{h}_{d}(t)$ denotes the dominant hands’ joint states, it is not impacted by the change of coordinates.
+Similarly, we apply the same modifications to the asymmetry-only action $a_{asym}(t)$, yielding $a_{\mathrm{AsymDex}}(t)=(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t))$, where $\hat{\xi}^{b}_{r}(t)$ is the target relative pose of the dominant hand now defined relative to the object being held by the facilitating hand.
+Incorporating the above change of coordinates in addition to leveraging asymmetry, allows us to define AsymDex’s policy as $\pi_{\mathrm{AsymDex}}(a_{\mathrm{AsymDex}}(t)|z_{\mathrm{AsymDex}}(t))$.
+Note that our formulation has significantly reduced the dimensions of both the state and action spaces, compared to the symmetric policy $\pi_{\mathrm{sym}}$ as defined in Section 3.
+We parameterize the AsymDex policy $\pi_{\mathrm{AsymDex}}(\cdot)$ using an MLP and using Proximal Policy Optimization (PPO) [49] to train it.
+See Appendix.
+A for details of the algorithm and policy architecture.
+4.3 Relative Pose Tracking Controller
+To control the hand bases based on the target relative pose $\hat{\xi}^{b}_{r}(t)$ provided by $\pi_{\mathrm{AsymDex}}$, we designed a bimanual controller that computes both the target dominant hand base pose $\hat{\xi}^{b}_{d}(t)$ and the target facilitating hand base pose $\hat{\xi}^{b}_{f}(t)$ as follows
+$\begin{split}\hat{\xi}^{b}_{d}(t)&=\alpha R^{o_{f}}_{world}\cdot dist(\hat{\xi}^{b}_{r}(t),\xi^{b}_{r}(t))+\xi^{b}_{d}(t),\\ \hat{\xi}^{b}_{f}(t)&=(\alpha-1)R^{o_{f}}_{world}\cdot dist(\hat{\xi}^{b}_{r}(t),\xi^{b}_{r}(t))+\xi^{b}_{f}(t),\end{split}$ (1)
+where $R^{o_{f}}_{world}$ denotes the rotational transformation from Frame $P_{f}$ to the world frame $P_{W}$, $dist(\cdot)$ denotes the difference between two 6D poses, and $\alpha$ is a hyperparameter that controls the relative involvement of each hand.
+The pseudo-code of the training process is included in Alg.
+1.
+4.4 Acquisition Phase
+Figure 3: We decompose asymmetric bimanual dexterous manipulation into two phases: An acquisition phase and an interaction phase.
+We show that AsymDex can be readily integrated with learned grasping policies in order to seamlessly acquire and manipulate objects.
+While our approach as explained thus far deals with the challenge of coordinating two hands to accomplish asymmetric dexterous manipulation tasks, it assumes that the object(s) of interest have already been grasped at the beginning of the task.
+However, in practice, robots must be learn to grasp and pick up the necessary objects before the interaction between the two hands (and the objects) can begin.
+We refer to this initial phase as the acquisition phase.
+Most recent works on bimanual dexterous manipulation often entirely ignore the acquisition phase and focus purely on the interaction phase [6, 19, 20].
+In contrast, we demonstrate that our approach can seamlessly accommodate the acquisition phase by i) leveraging the observation that the acquisition phase doesn’t require the coordination of two arms, and ii) employing recent advances in learning to grasp.
+Specifically, we demonstrate that we can seamlessly integrate AsymDex with PDGM [50], which can efficiently learn multi-fingered grasping policies by leveraging pre-grasp poses (see Fig.
+3).
+Details about the grasping reward design are available in Appendix.
+B.
+We begin by executing the grasping policy in isolation and then ”turn on” the asymmetric policy learned by AsymDex after the object has been firmly grasped by the facilitating hand.
+If the task requires the dominant hand to also grasp a second object, we employ the same method to train a grasping policy for dominant hand to acquire the object, but switch the control of the dominant hand’s joints over the asymmetric policy after the object has been grasped.
+5 Experimental evaluation
+We evaluated AsymDex on four simulated and three real-world bimanual dexterous tasks and compared its performance against strong baselines that challenge our key design choices.
+5.1 Simulation Experiments
+Our experiments in simulation both systematically and rigorously evaluate AsymDex.
+Tasks: We evaluated AsymDex and the baselines on the following four bimanual manipulation tasks which contain both original (Block in cup, Bottle cap) and adapted tasks (Stack, Switch) from BiDexHand [14] (see Fig.
+1).
+All these tasks use two Shadow Hands – each a 30-DoF simulated multi-finger hand system (24-DoF hand + 6-DoF floating wrist base) built with Isaac Gym [51].
+  • $\bullet$
+    Block in cup: The two hands must coordinate to ensure that one hand places a block inside a cup that is being held by the other without letting either the cup or the block fall to the ground.
+  • $\bullet$
+    Stack: Two cups need to be stacked together. Each hand must hold a cup, and both must coordinate such that the two cups are aligned as one slides into the other.
+  • $\bullet$
+    Bottle cap: One hand must hold and reorient a bottle such that the other hand can grasp and separate the bottle cap from the bottle.
+  • $\bullet$
+    Switch: One hand holds and reorients a switch in a way that allows the other hand to turn it on.
+Note that our task designs are more challenging than their counterparts in BiDexHand [14].
+We require that the two hands coordinate and synchronize to achieve success in each of the above four tasks, especially since (unlike the original designs) we do not provide a support surface (e.g., a table) that would significantly reduce the need for bimanual coordination.
+See Appendix.
+C for details on state space design, sampling procedure, success criteria, and reward design.
+Metrics: We quantify performance in terms of i) success rate (see Appendix. C for criteria) and ii) sample efficiency. We report both metrics across five random seeds in all experiments.
+5.1.1 Learning Bimanual Coordination
+We first evaluated AsymDex’s effectiveness during the interaction phase.
+Following contemporary practice in methods that learn bimanual dexterous skills [19, 20], we initialized the environment such that the hands are at appropriate pre-grasp poses near the respective objects.
+This allows us to isolate and examine AsymDex’s ability to learn to coordinate two multi-fingered hands.
+See Section 5.1.2 for the second experiment in which we also consider the challenge of acquiring the objects from a tabletop surface before interaction begins.
+We compared AsymDex against the following baselines:
+  • $\bullet$
+    Sym: This policy assumes that both hands play an equal role in bimanual manipulation (see $\pi_{sym}$ in Sec.
+3).
+This baseline allows us to examine the necessity and effectiveness of leveraging the asymmetry in hand roles as well as the relative action and observation spaces.
+  • $\bullet$
+    Asym-w/o-rel: This policy leverages asymmetry in hand roles, but learns over absolute observation and action places (see Sec.
+4.1).
+As such, this baseline allows us to examine the necessity and effectiveness of relative action and observation spaces.
+  • $\bullet$
+    Rel-w/o-asym: This policy leverages the relative observation and action places, but ignores asymmetry. As such, it allows us to examine the necessity and effectiveness of asymmetry.
+Figure 4: AsymDex consistently outperforms the baselines in terms of sample efficiency and success rate. Solid lines indicate mean trends and shaded areas show $\pm$ std. dev., over five random seeds.
+Both asymmetry and relative spaces are necessary for consistent performance: We report the learning curves in Fig.
+4 (a) and success rates in Table 1.
+Note that AsymDex consistently outperforms all the baselines across all four tasks in terms of success rate and sample efficiency, with significant margins in two tasks (Block in cup and Switch).
+Rel-w/o-asym performs better than the other two baselines across all tasks except Switch, while Asym-w/o-rel performs better than Monolithic on all tasks except Block in cup, on which both struggle.
+Taken together, the above observations reveal a few key insights.
+First, when used in isolation, neither asymmetry nor relative spaces are sufficient across all tasks.
+Second, the use of relative spaces offers a larger boost in performance compared to asymmetry, likely due to the fact that relative spaces avoid unnecessary exploration (e.g., when the two hands move in parallel) while allowing the facilitating hand to exhibit more complex behaviors.
+Third, ignoring both asymmetry and relative spaces (Sym) hardly leads to success.
+Task Method Sym Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
+Block in cup $0.0429\pm 0.0266$ $0.0164\pm 0.0190$ $0.1086\pm 0.1378$ 0.7701 $\pm$ 0.0559
+Stack $0.0771\pm 0.0611$ $0.2185\pm 0.3232$ $0.6560\pm 0.3213$ 0.8392 $\pm$ 0.0596
+Bottle cap $0.3111\pm 0.1813$ $0.4143\pm 0.1410$ $0.4730\pm 0.2011$ 0.6295 $\pm$ 0.1422
+Switch $0.0563\pm 0.0126$ $0.1626\pm 0.0882$ $0.1149\pm 0.0176$ 0.6700 $\pm$ 0.0359
+Table 1: Success rates (mean $\pm$ std. dev.) for the interaction phase.
+5.1.2 Learning to Grasp and Coordinate
+We next evaluated AsymDex’s ability to incorporate the object acquisition phase before the interaction phase.
+Specifically, we initialize the environment for each task such that the objects of interest are placed on a tabletop surface.
+As such, each method needs to learn both to grasp the necessary objects and to coordinate the two hands to complete the tasks.
+Table 2: Success rates (mean $\pm$ std. dev.) after combining acquisition and interaction phases
+Task Method Monolithic 2-stage-sym AsymDex
+Block in cup $0.0321\pm 0.0251$ $0.1505\pm 0.1059$ 0.7938 $\pm$ 0.0897
+Bottle cap $0.0\pm 0$ 0.0 $0.2552\pm 0.1573$ 0.6116 $\pm$ 0.1328
+For AsymDex, we follow the same strategy introduced in Section 4.4, and comapare its performance against the following baselines:
+  • $\bullet$
+    Monolithic: This baseline uses a single policy to learn both the grasping and interaction phases for both hands, allowing us to investigate the benefits of two-phase decomposition.
+  • $\bullet$
+    2-stage-sym: This policy benefits from the two phase decomposition but leverages neither asymmetry nor relative spaces.
+As such, this baseline allows us to examine if merely employing two-phase decomposition is sufficient.
+To ensure a fair comparison, we provide pre-grasp pose annotations to both baselines.
+Further, we ensure that the total number of env.
+interactions (the number of one stage or the sum of two stages) is the same across AsymDex and the baselines.
+See Appendix.
+B for details of the grasping learning.
+AsymDex can effectively combine the acquisition and interaction phases: We report the overall roll-out success rate of all methods for two tasks across five random seeds in Table.
+2.
+We find that AsymDex significantly outperforms the other two baselines in both tasks, suggesting that combining phase decomposition with AsymDex’s other two design choices (asymmetry and relative spaces) results in policies that can effectively handle both the acquisition and the interaction phases of bimanual dexterous manipulation.
+The fact that 2-stage-sym baseline outperforms the Monolithic baseline points to the inherent benefits of phase decomposition.
+Our qualitative analysis of Block in cup task revealed that Monolithic policy learns to tip the cup over and push the block towards the cup.
+In contrast, both the two-stage policies learn more intuitive behaviors, suggesting that the phase decomposition nudges the grasping and interaction policies to learn reasonable behaviors that complement each other.
+5.2 Real-world Experiments
+We finally evaluated AsymDex and the same baselines on the following 3 real-world bimanual manipulation tasks inspired by recent bimanual manipulation works [5, 19, 26, 52, 53, 54, 55].
+Hardware setup: We use one 16-dof Allegro hand from Wonik Robotics and one 6-DoF Ability Hand from Psyonic.
+They are mounted on two 7-DoF Kinova Gen3 robotic arms.
+We used the Ability Hand for one of the hands since we only had access to one physical Allegro hand.
+For object pose information, we either employ an Realsense camera with AprilTag-based object tracking method or estimate it directly from the end-effector’s pose, especially during occlusion.
+Real-world Tasks: i) Block in Cup: Place a block with sides of 5 cm inside a cylindrical cup with an internal diameter of 10 cm, ii) Pour: Pour dry beans from one cup into another, and iii) Twist Lid: Twist the lid off a jar while holding the jar.
+Snapshots are shown in Fig.
+5.
+Simulation Evaluation: For each task, we first construct a simulated counterpart of our hardware setup and then train all policies in simulation.
+We then compared their performance in terms of success rate and sample efficiency, as reported in Fig.
+4 (b).
+As observed in previous experiments, AsymDex consistently and significantly outperformed all baselines in terms of both metrics.
+Sim2Real Transfer: We deployed only the AsymDex policies on hardware since the baseline policies tended to exhibit either negligible success rates or aggressive behaviors that could damage the hardware.
+To enable effective sim-to-real transfer, we apply domain randomization during reinforcement learning, and details of the randomization process are provided in Appendix.
+D.
+Block in cup Pouring Twist Lid
+16/20 17/20 18/20
+Table 3: Real-world task success rates for AsymDex.
+AsymDex enables zero-shot Sim2Real Transfer: As shown in Table.
+3, AsymDex achieves high success rates across all deployments.
+This once again highlights the effectiveness of AsymDex in learning robust and reliable real-world bimanual dexterous manipulation skills.
+For the block in cup and pour task, AsymDex is simply trained with observation noise component of common domain randomization technique to learn a policy that can be deployed in real world successfully.
+For the most challenging twist lid task, we added observation noise and action noise, as well as other randomization components for sim2real transfer.
+Though we utilized similar reward design as previous work [19] for this task, we observed that AsymDex learns a more natural and human-like behavior compared to the previous work.
+Besides, unlike previous work which fixes the hand base motion [2, 19], AsymDex is able to reposition and reorient both hand bases while two hands interact, enabling more fluent bimanual coordination.
+See Appendix.
+D for the detailed reward design and simulation success criterion for each task.
+Figure 5: We created simulation environments to match our hardware setup.
+6 Conclusion
+Our framework (AsymDex) is capable of learning complex asymmetric bimanual dexterous manipulation tasks via reinforcement without relying on demonstrations.
+We introduced and validated the need for AsymDex’s two crucial ingredients: assigning asymmetric roles to the two hands, and using relative observation and action spaces.
+Our evaluation results reveal that the combination of these choices consistently leads to better sample efficiency and success rates across different tasks.
+7 Limitations and Future Work
+Our work has revealed a number of limitations and avenues for future research.
+First, AsymDex in its current form cannot handle certain bimanual tasks that require complex multi-finger manipulation from both hands (e.g., reorienting a heavy object, dynamic handover).
+Second, AsymDex does not consider the kinodynamic constraints that might result from manipulator arms.
+Three, behaviors produced by AsymDex are not always natural or human-like due to lack of necessary incentives.
+Appendices
+Appendix A Training Details
+We use Proximal Policy Optimization (PPO) [49] algorithm to train all policies $\pi_{\mathrm{naive}},\pi_{\mathrm{asym}},$ and $\pi_{\mathrm{AsymDex}}$ with their corresponding value functions.
+Both policies and value functions are parameterized via a three-layer MLP network.
+The size of hidden layers for each is i) policy: (256, 256, 128), ii) value function: (512, 512, 512).
+The activation functions are all set as Exponential Linear Unit (ELU).
+We use the same PPO hyperparameters for all the baselines and AsymDex ($\gamma:0.98,\lambda:0.95$, clip range: 0.2, minibatch size: 8092).
+We use an adaptive learning rate with KL threshold of 0.016.
+We train the polices on a computer with a single Nividia RTX 4090 GPU.
 1 Randomly initialize the two hand bases’ poses $\xi^{b}_{f}(0)$ and $\xi^{b}_{d}(0)$, and initialize object poses $o_{f}(0)$ and $o_{d}(0)$ based on $\xi^{b}_{f}(0)$ and $\xi^{b}_{d}(0)$.
@@
 Algorithm 1 AsymDex: Interaction Phase
-In addition to asymmetry, a key characteristic of bimanual dexterous manipulation is the synchronized motion of the two hands in which each hand moves in response to the other.
-In this section, we explain how we can further reduce the size of the observation and action spaces by defining relative and object-centric spaces that capture the relationships between the motions of two hands and the object(s) being manipulated.
-Indeed, the use of relative state spaces has shown to considerably benefit bimanual manipulation with simple end effectors [42, 43, 44, 23].
-Furthermore, some of these prior work is limiting the relative space in a one-degree-of-freedom (1-DoF) action space [23], while AsymDex allows for complete 6-DoFs relative space.
-Let $o_{f}=o_{1}$ be the state of the object being held by the facilitating hand, and let $o_{d}=o_{2}$ be the state of the object being manipulated by the dominant hand.
-We attach a coordinate frame to the object being held by the facilitating hand: $P_{f}$.
-Now, we can transform the observations $(\xi^{b}_{d}(t),\xi^{b}_{f}(t),\xi^{h}_{d}(t),o_{f}(t),o_{d}(t))$, which where originally defined in the world coordinate frame $P_{W}$, into the new coordinate frame $P_{f}$.
-Note that since there is no relative motion between the facilitating hand and the object that it’s holding, neither $o_{f}(t)$ nor $\xi^{b}_{f}(t)$ do not change in $P_{f}$.
-As such, both $o_{f}(t)$ and $\xi^{b}_{f}(t)$ can be neglected without losing any information.
-Now, transforming the remaining observations $(\xi^{b}_{d}(t),\xi^{h}_{d}(t),o_{d}(t))$ into the new coordinate $P_{f}$ yields $(\xi^{b}_{r}(t),\xi^{h}_{d}(t),o_{r}(t))$, respectively.
-Here, $\xi^{b}_{r}(t)$ denotes the 6D pose of the dominant hand base and $o_{r}(t)$ denotes the 6D pose of the object being manipulated by the dominant hand, both now defined relative to the object being held by the facilitating hand.
-Note that since $\xi^{h}_{d}(t)$ denotes the dominant hands’ joint states, it is not impacted by the change of coordinates.
-Similarly, we apply the same modifications to the action spaces of the asymmetry bimanual policy.
-This allows us to reduce the actions from $(\hat{\xi}^{b}_{d}(t),\hat{\xi}^{b}_{f}(t),\hat{\xi}^{h}_{d}(t))$ to $(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t))$, where $\hat{\xi}^{b}_{r}(t)$ is the target relative pose of the dominant hand now defined relative to the object being held by the facilitating hand.
-Incorporating the above change of coordinates in addition to leveraging asymmetry, allows us to define AsymDex’s policy as $\pi_{\mathrm{AsymDex}}(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t)|\xi^{b}_{r}(t),\xi^{h}_{d}(t),o_{r}(t))$.
-Note that our formulation has significantly reduced the dimensions of both the state and action spaces, compared to the naive policy $\pi_{\mathrm{naive}}$ as defined in Section 3.2.
-See Appendix.
-A for RL algorithm and policy architecture.
-3.5 Relative Pose Controller
-To control the hand bases based on the target relative pose $\hat{\xi}^{b}_{r}(t)$ provided by $\pi_{\mathrm{AsymDex}}$, we designed a bimanual controller that computes both the target dominant hand base pose $\hat{\xi}^{b}_{d}(t)$ and the target facilitating hand base pose $\hat{\xi}^{b}_{f}(t)$ as follows
-$\begin{split}\hat{\xi}^{b}_{d}(t)&=\alpha R^{o_{f}}_{world}\cdot dist(\hat{\xi}^{b}_{r}(t),\xi^{b}_{r}(t))+\xi^{b}_{d}(t),\\ \hat{\xi}^{b}_{f}(t)&=(\alpha-1)R^{o_{f}}_{world}\cdot dist(\hat{\xi}^{b}_{r}(t),\xi^{b}_{r}(t))+\xi^{b}_{f}(t),\end{split}$ (1)
-where $R^{o_{f}}_{world}$ denotes the rotational transformation from Frame $P_{f}$ to the world frame $P_{W}$, $dist(\cdot)$ denotes the difference between two 6D poses.
-and $\alpha$ is a hyperparameter that controls the involvement of each hand.
-Throughout our experiments, we used fixed $\alpha=0.5$.
-The pseudo-code of the training process is included in Alg.
-1.
-3.6 Acquiring Objects
-Figure 2: We decompose asymmetric bimanual dexterous manipulation into two phases: An acquisition phase and an interaction phase.
-We show that AsymDex can be integrated with learned grasping policies in order to seamlessly execute both phases.
-While our approach as explained thus far deals with the challenge of coordinating two hands to accomplish asymmetric dexterous manipulation tasks, it assumes the task begins with that the object(s) of interest have already been grasped.
-However, in practice, robots must be able to tackle the challenge of grasping the necessary objects before the interaction between the two hands and the objects can begin.
-We refer to this phase as the acquisition phase.
-Most recent works on bimanual dexterous manipulation often entirely ignore the acquisition phase and focus purely on the interaction phase [7, 5, 40].
-In contrast, we demonstrate that our approach can seamlessly accommodate the acquisition phase by i) leveraging the observation that the acquisition phase doesn’t require the coordination of two arms, and ii) employing recent advances in learning to grasp.
-Specifically, we demonstrate that we can seamlessly integrate AsymDex with PDGM [45], which can efficiently learn multi-fingered grasping policies by leveraging pre-grasp poses (see Fig.
-2).
-Details about the grasping reward design are available in Appendix.
-B.
-We begin by executing the grasping policy in isolation and then ”turn on” the asymmetric policy learned by AsymDex after the object has been firmly grasped by the facilitating hand.
-If the task requires the dominant hand to also grasp a second object, we employ the same method to train a grasping policy for dominant hand to acquire the object, but switch the control of the dominant hand’s joints over the asymmetric policy after the object has been grasped.
-4 Evaluation
-We evaluated AsymDex on four asymmetric bimanual manipulation tasks and compared its performance against strong baselines that challenge our key design choices.
-We begin by explaining the various aspects of our experimental setup and follow with a discussion of results.
-Platform: We conducted all our experiments using two ShadowHands – each a 30-DoF simulated multi-finger hand system (24-DoF hand + 6-DoF floating wrist base) built with Isaac Gym [46].
-Tasks: We evaluated AsymDex and the baselines on the following four biannual manipulation tasks which contain both original (Block in cup, Bottle cap) and adapted tasks (Stack, Switch) from BiDexHand [20] (see Fig.
-3 for visualization).
-  • $\bullet$
-    Block in cup: The two hands must coordinate to ensure that one hand places a block inside a cup that is being held by the other without letting either the cup or the block fall to the ground.
-  • $\bullet$
-    Stack: Two cups need to be stacked together. Each hand must hold a cup, and both must coordinate such that the two cups are aligned as one slides into the other.
-  • $\bullet$
-    Bottle cap: One hand must hold and reorient a bottle such that the other hand can grasp and separate the bottle cap from the bottle.
-  • $\bullet$
-    Switch: One hand holds and reorients a switch in a way that allows the other hand to turn it on.
-Note that the hands have to coordinate and perform all four tasks without relying on a support surface such as a table and ensure that the object(s) are not dropped.
-See Appendix.
-C for details on state space design, sampling procedure, success criteria, and reward design for each task.
-Figure 3: We evaluate AsymDex on four bimanual dexterous manipulation tasks.
-Metrics: We quantify performance in terms of i) Task success rate (see Appendix. C for criteria) and ii) sample efficiency. We report both metrics across five random seeds in all experiments.
-We evaluated AsymDex with two sets of experiments as described below.
-4.1 Learning Bimanual Coordination
-In this experiment, we focus on AsymDex’s effectiveness during the interaction phase.
-We initialize the environment such that the hands are at a pre-grasp pose around the objects using appropriate initial grasps (After the first timestep of the environment, the hand need to learn to catch and grasp the object from pre-grasp pose if it is not a facilitating hand).
-Note that this is a common assumption in recent methods that learn bimanual dexterous manipulation skills [5, 40].
-Further, this allows us to isolate and examine AsymDex’s ability to learn to coordinate two multi-fingered hands.
-See Section 4.2 for the second experiment in which we also consider the challenge of acquiring the objects from a tabletop surface before interaction begins.
-We compare AsymDex policy against the following baselines:
-  • $\bullet$
-    Monolithic: This policy doesn’t make assumptions about the structure of bimanual manipulation (see Sec.
-3.2).
-As such, this baseline allows us to examine the necessity and effectiveness of leveraging both the asymmetry in hand roles and the relative action and observation spaces.
-  • $\bullet$
-    Asym-w/o-rel: This policy leverages asymmetry in hand roles, but learns over absolute observation and action places (see Sec.
-3.3).
-As such, this baseline allows us to examine the necessity and effectiveness of relative action and observation spaces.
-  • $\bullet$
-    Rel-w/o-asym: This policy leverages the relative observation and action places, but ignores asymmetry. As such, it allows us to examine the necessity and effectiveness of asymmetry.
-Figure 4: AsymDex consistently outperforms the baselines in terms of sample efficiency and success rate. Solid lines indicate mean trends and shaded areas show $\pm$ std. dev., over five random seeds.
-We report the learning curves in Fig.
-4 and success rates in Table 1.
-As can be readily observed, AsymDex is the only method that consistently performs well across all tasks, either performing comparably or outperforming all the baselines.
-While Rel-w/o-asym performs comparably to AsymDex on Stack, it is not able to match AsymDex’s performance on the other tasks.
-Both AsymDex and Rel-w/o-asym perform substantially better than the other two baselines across all tasks except Switch, where AsymDex performs much better than all three baselines.
-In fact, Asym-w/o-rel is able to learn an effective policy in only one task (Bottle cap), while Monolithic struggles on all four tasks.
-Curiously, Monolithic outperforms asym-w/o-rel on the Stack task.
-Qualitative analysis of rollouts reveals that, unlike Asym-w/o-rel, the Monolithic policy learns to use the facilitating hand’s fingers to orient the cup towards the bottom of the other cup.
-Table 1: Success rates (mean $\pm$ std. dev.) in Interaction phase
-Task Method Monolithic Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
-Block in cup $0.0429\pm 0.0266$ $0.0164\pm 0.0190$ $0.3052\pm 0.1577$ 0.7701 $\pm$ 0.0559
-Stack $0.2517\pm 0.2809$ $0.0044\pm 0.0037$ 0.9443 $\pm$ 0.0136 $0.9233\pm 0.0250$
-Bottle cap $0.6125\pm 0.1734$ $0.7893\pm 0.1277$ $0.8073\pm 0.1389$ 0.8301 $\pm$ 0.1797
-Switch $0.0563\pm 0.0126$ $0.1626\pm 0.0882$ $0.1149\pm 0.0176$ 0.6700 $\pm$ 0.0359
-Taken together, the above observations reveal a few insights.
-First, when used in isolation, neither asymmetry nor relative motion are sufficient across all tasks.
-In particular, while each might provide sufficient structure to accomplish some tasks, they prove to be less effective on other tasks.
-Second, the use of relative motion offers a larger boost in performance compared to asymmetry, likely due to the fact that relative spaces avoid unnecessary exploration (e.g., when the two hands move in parallel) while allowing the facilitating hand to exhibit more complex behaviors.
-Third, ignoring both asymmetry and relative motion hardly leads to success.
-4.2 Learning to Grasp and Coordinate
-In this experiment, we evaluate AsymDex’s ability to incorporate the object acquisition phase in addition to the interaction phase.
-Specifically, we initialize the environment for each task such that the objects of interest are placed on a tabletop surface.
-As such, each method needs to learn both to grasp the necessary objects and to coordinate the two hands to complete the tasks.
-For AsymDex, we follow the same strategy as introduced in Section 3.6. We compare AsymDex’s performance against the following baselines:
-  • $\bullet$
-    1-stage-monolithic: This baseline uses a single policy to learn both the grasping and interaction phases for both hands, and thus allows us to investigate the benefits of AsymDex’s two-phase decomposition.
-  • $\bullet$
-    2-stage-monolithic: This policy benefits from the same phase decomposition as AsymDex policy, but leverages neither asymmetry nor relative motion.
-As such, this baseline allows us to examine if this task can be solved merely with two-phase decomposition.
-To ensure a fair comparison, we provide pre-grasp pose annotations to both baselines.
-Further, we ensure that the total number of env.
-interactions (the number of one stage or the sum of two stages) is the same across AsymDex and the baselines.
-See Appendix.
-B for details of the grasping learning.
-We report the overall roll-out success rate of all methods for two tasks across five random seeds in Table.
-2.
-We find that AsymDex significantly outperforms the other two baselines in both tasks, suggesting that combining phase decomposition with AsymDex’s other two design choices (asymmetry and relative state) results in policies that can effectively handle both the acquisition and the interaction phases of bimanual dexterous manipulation.
-The fact that 2-stage-monolithic baseline outperforms the 1-stage-monolithic baseline demonstrates the inherent benefits of phase decomposition.
-Our qualitative analysis of Block in cup task revealed that 1-stage-monolithic policy learns to tip the cup over and push the block towards the cup.
-In contrast, both the two-stage policies learn the expected behavior.
-This suggests that the phase decomposition nudges the grasping and interaction policies to learn reasonable behaviors that complement each other.
-Table 2: Success rates (mean $\pm$ std. dev.) after combining acquisition and interaction phases
-Task Method 1-stage-monolithic 2-stage-monolithic 2-stage-AsymDex
-Block in cup $0.0321\pm 0.0251$ $0.1505\pm 0.1059$ 0.7938 $\pm$ 0.0897
-Bottle cap $0.1680\pm 0.2695$ $0.6407\pm 0.1141$ 0.8726 $\pm$ 0.0600
-5 Conclusion
-Our framework (AsymDex) is capable of learning complex asymmetric bimanual dexterous manipulation tasks via reinforcement without relying on demonstrations.
-We introduced and validated the need for AsymDex’s two crucial ingredients: assigning asymmetric roles to the two hands, and using relative observation and action spaces.
-Our evaluation results reveal that the combination of these choices consistently leads to better sample efficiency and success rates across different tasks.
-6 Limitations and Future Works
-Our work has revealed a number of limitations and avenues for future research.
-First, AsymDex in its current form cannot handle certain bimanual tasks that require complex multi-finger manipulation from both hands (e.g., reorienting a heavy object, dynamic handover).
-Second, AsymDex does not consider the kinodynamic constraints that might result from manipulator arms.
-Third, AsymDex has not yet been evaluated on hardware.
-Fourth, behaviors produced by AsymDex are not always natural or human-like due to lack of necessary incentives.
-Appendices
-Appendix A RL Training
-We use Proximal Policy Optimization (PPO) [47] algorithm to train all policies $\pi_{\mathrm{naive}},\pi_{\mathrm{asym}},$ and $\pi_{\mathrm{AsymDex}}$ with their corresponding value functions.
-Both policies and value functions are parameterized via a three-layer MLP network.
-The size of hidden layers for each is i) policy: (256, 256, 128), ii) value function: (512, 512, 512).
-The activation functions are all set as Exponential Linear Unit (ELU).
-We use the same PPO hyperparameters for all the baselines and AsymDex ($learning\ rate:3\times 10^{-4},\gamma:0.98,\lambda:0.95$).
-We train the polices on a computer with a single Nividia RTX 4090 GPU.
 Appendix B Grasping Learning
 Two-stage policy
-Both AsymDex (our approach) and 2-stage-monolithic policy (one of the baselines in Sec.
-4.2) are two-stage policies.
+Both AsymDex (our approach) and 2-stage-sym policy (one of the baselines in Sec.
+5.1.2) are two-stage policies.
 Therefore, they can first learn a grasping policy for the facilitating hand (or two grasping policies for facilitating hand and dominant hand respectively).
-Such policy $\pi_{\mathrm{grasp}}(\hat{\xi}^{h}_{f}(t)|\xi^{h}_{f}(t),\xi^{b}_{f}(t)\circleddash X_{o_{f}})$ takes in the hand joint states and the relative pose between hand and the object, and outputs the target hand joint positions to grasp the object firmly.
+Such policy $\pi_{\mathrm{grasp}}(\hat{\xi}^{h}_{f}(t)|\xi^{h}_{f}(t),\xi^{b}_{f}(t)\circleddash P_{f})$ takes in the hand joint states and the relative pose between hand and the object, and outputs the target hand joint positions to grasp the object firmly.
 We first provide pre-grasp annotations [50], which allows the hands to initialize at the position close to the objects with proper joint positions.
@@
 Another baseline in Sec.
-4.2, i.e., the 1-stage-monolithic policy, does not incorporate the task decomposition.
+5.1.2, i.e., the monolithic policy, does not incorporate the task decomposition.
 Therefore, it only uses the task-specific interaction rewards (see Appendix.
@@
 For a fair comparison, both hands also start at the pre-grasp poses.
-Appendix C Task Design
-In this section, we show the details for each task.
+Appendix C BiDexHands Task Design
+In this section, we show the details for each BiDexHands simulation task.
 State Space Design
@@
 And for all policies, we also include the previous actions in the policy input.
-For the block in cup task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the cup and the block respectively.
-For the stack task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of two cups.
-For the Bottle cap task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the bottle and the cap respectively.
-For the Switch task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the switch body and the button respectively.
+For the block in cup task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the cup and the block respectively.
+For the stack task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of two cups.
+For the Bottle cap task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the bottle and the cap respectively.
+For the Switch task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the switch body and the button respectively.
 The dimensions of the observation and action spaces of each policy are shown in Table.
-3.
+4.
 It is obvious that AsymDex policy significantly reduces the state dimensions.
 Table 4: Dimension of Observation and action spaces. For all tasks, the dimensions are identical.
-Monolithic Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
+Sym Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
 Observation $176$ $108$ $163$ $88$
@@
 Sampling Procedure
-  • Place block in cup: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$.
+  • Block in cup: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$.
 For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.57,1.57)$, in radians.
@@
 The initial position of the facilitating hand base is at $[0.55,0.6,0.8]$.
-  • Stack cups: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$.
+  • Stack: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$.
 For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.57,1.57)$, in radians.
@@
 The initial position of the facilitating hand base is at $[0.55,0.6,0.8]$.
-  • Open bottle cap: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.58,0.62)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.21,-0.19)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.58,0.62)$.
+  • Bottle cap: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.58,0.62)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.21,-0.19)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.58,0.62)$.
 For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.0,1.0)$, in radians.
@@
 Thus its position and rotation is calculated based on the initial position and rotation of facilitating hand base.
-  • Turn on switch: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.2,0.6)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.25,-0.05)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.5,0.9)$.
+  • Switch: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.2,0.6)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.25,-0.05)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.5,0.9)$.
 For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.0,1.0)$, in radians.
@@
 Success Criteria
-  • Place block in cup The task is considered successful if the distance of the block center and the cup center is smaller than 0.035 meters.
+  • Block in cup The task is considered successful if the distance of the block center and the cup center is smaller than 0.035 meters.
 This distance makes sure the task is only considered successful when the block is inside the cup.
 If the block falls on the ground or has not entered the cup within a certain time step, the task is considered failed.
-  • Stack cups The task is considered successful if the distance between the cup centers is smaller than 0.02 meters.
+  • Stack The task is considered successful if the distance between the cup centers is smaller than 0.02 meters.
 If either cup falls on the ground or has not been stacked within a certain time step, the task is considered failed.
-  • Open bottle cap The task is considered successful if the cap is taken off from its original position 0.05 meters away within a time duration, and is considered failed otherwise.
-  • Turn on switch The button and the switch body are connected by a revolute joint ranging from 0 to 0.5585 rads.
+  • Bottle cap The task is considered successful if the cap is taken off from its original position 0.05 meters away within a time duration, and is considered failed otherwise.
+  • Switch The button and the switch body are connected by a revolute joint ranging from 0 to 0.5585 rads.
 The task is considered successful if the button is pressed and rotated 0.3585 rads within a time duration, and is considered failed otherwise.
@@
 The reward design is similar across all tasks:
-$Reward=R_{hand\ distance}+R_{progress}+R_{action\ penalty}+R_{success\ bonus}$
+$Reward=\alpha_{1}R_{hand\ distance}+\alpha_{2}R_{progress}+\alpha_{3}R_{action\ penalty}+\alpha_{4}R_{success\ bonus}$
 For each task, $R_{action\ penalty}=-||a(t)||^{2}$, and the $R_{success\ bonus}$ is the task success reward. $R_{hand\ distance}$ and $R_{progress}$ are slightly different for each tasks.
-  • Place block in cup: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth.
+  • Block in cup: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth.
 $R_{progress}=-||x_{cup}-x_{block}||$, where $x_{cup}$ is the position of the cup, and $x_{block}$ is the position of the block.
-  • Stack cups: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth, which is grasped by the facilitating hand.
+  • Stack: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth, which is grasped by the facilitating hand.
 $R_{progress}=-||x_{cupd}-x_{cupf}||$, where $x_{cupf}$ is the position of the cup grasped by the facilitating hand, and $x_{cupd}$ is the position of the cup grasped by the dominant hand.
-  • Open bottle cap: $R_{hand\ distance}=(1-(||x_{index}-x_{cap}||+||x_{thumb}-x_{cap}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{cap}$ is the position of the bottle cap.
+  • Bottle cap: $R_{hand\ distance}=(1-(||x_{index}-x_{cap}||+||x_{thumb}-x_{cap}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{cap}$ is the position of the bottle cap.
 $R_{progress}=||x_{cap}-x_{bottle\ top}||$, where $x_{cap}$ is the position of the cap, and $x_{bottle\ top}$ is the position of the top of the bottle.
-  • Turn on switch: $R_{hand\ distance}=(1-(||x_{index}-x_{button}||+||x_{thumb}-x_{button}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{button}$ is the position of the button.
+  • Switch: $R_{hand\ distance}=(1-(||x_{index}-x_{button}||+||x_{thumb}-x_{button}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{button}$ is the position of the button.
 $R_{progress}=2*\theta_{button}$, where $\theta_{button}$ is the rotated angle of the joint that connects the button and the switch body.
+Appendix D Real-world Task Design
+In this section, we present the details for each real-world task.
+State Space Design
+For each task, the hand joint states $\xi^{h}_{f}(t)$, $\xi^{h}_{d}(t)$ include the 16-DoF hand joint positions.
+We use quaternions to represent rotation part of object and hand base poses.
+For the block in cup task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the cup and the block respectively.
+For the pour task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of two cups.
+For the twist lid task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the jar and the lid respectively.
+The dimensions of the observation and action spaces of each policy are shown in Table.
+5.
+AsymDex policy significantly reduces the state dimensions in real-world tasks consistently.
+Table 5: Dimension of Observation and action spaces. For all tasks, the dimensions are identical.
+Sym Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
+Observation $60$ $44$ $53$ $30$
+Action $44$ $28$ $38$ $22$
+Sampling Procedure
+The corresponding simulation environments of real-world tasks include two arms and two hands attached to them.
+Hence, we randomize the initial position of objects and hand poses by randomizing the initial joint angles of two Kinova arms.
+Then we initialize the poses of objects according to the initial hand poses.
+  • Block in cup: The default initial joint angles of the 7-DoF Kinova arms attached to the dominant hand are $[0.0,0.8,0,\pi/2+0.5,0,-1.3,-\pi/2]$.
+We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.1,0.1),\quad\forall i\in\{1,\dots,4\}$; $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.2,0.2),\quad\forall i\in\{5,6\}$; $\delta\theta_{7}\in\Theta_{7}\sim\mathcal{U}(-0.3,0.3)$.
+The default initial joint angles of the 7-DoF Kinova arms attached to the facilitating hand are $[-0.0,0.8,0,\pi/2+0.5,0,-1.3,0]$.
+We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.0,0.0),\quad\forall i\in\{1,\dots,3\}$; $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.1,0.1),\quad\forall i\in\{4,7\}$.
+  • Pour: The default initial joint angles of the 7-DoF Kinova arms attached to the dominant hand are $[0.2,0.8,0,\pi/2+0.5,0,-1.3,-\pi/2]$.
+We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.1,0.1),\quad\forall i\in\{1,\dots,4\}$; $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.2,0.2),\quad\forall i\in\{5,6\}$; $\delta\theta_{7}\in\Theta_{7}\sim\mathcal{U}(-1.1,0.6)$.
+The default initial joint angles of the 7-DoF Kinova arms attached to the facilitating hand are $[-0.2,0.8,0,\pi/2+0.5,0,-1.3,0]$.
+We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.0,0.0),\quad\forall i\in\{1,\dots,3\}$; $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.1,0.1),\quad\forall i\in\{4,7\}$.
+  • Twist lid: The default initial joint angles of the 7-DoF Kinova arms attached to the dominant hand are $[0.1,0.5,0,\pi/2+0.55,0.75,-1.5,-\pi/2-0.2]$.
+We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.0,0.0),\quad\forall i\in\{1,\dots,4\}$; $\delta\theta_{5}\in\Theta_{5}\sim\mathcal{U}(-0.05,0.05)$; $\delta\theta_{6}\in\Theta_{6}\sim\mathcal{U}(-0.2,0.0)$; $\delta\theta_{7}\in\Theta_{7}\sim\mathcal{U}(-0.1,0.1)$.
+The default initial joint angles of the 7-DoF Kinova arms attached to the facilitating hand are $[-0.1,0.8,0,\pi/2+0.5,-0.7,-1.35,-0.1]$.
+Success Criteria
+  • Block in cup The task is considered successful if the distance of the block center and the cup center is smaller than 0.035 meters.
+This distance makes sure the task is only considered successful when the block is inside the cup.
+If the block falls on the ground or has not entered the cup within a certain time step, the task is considered failed.
+  • Pour The task is considered successful if the distance between the rims of the two cups is less than 0.035 meters and the cup grasped by the facilitating hand is up-right.
+If either cup falls on the ground or the task is not success within a certain time step, the task is considered failed.
+  • Twist lid We utilize the articulated bottle simulation of the previous work [19].
+The lid and bottle are connected by a single revolute joint.
+The task is considered successful if the lid is rotated over $3\times\pi$ rad within a time duration, and is considered failed otherwise.
+Reward Design
+The reward function structure is similar across all tasks:
+$Reward=R_{task}+\alpha_{2}R_{action\ penalty}+\alpha_{3}R_{success\ bonus}$
+For each task, $R_{action\ penalty}=-||a(t)||^{2}$, and the $R_{success\ bonus}$ is the task success bonus.
+  • Block in cup: $R_{task}=\beta_{1}R_{hand\ dist}+\beta_{2}R_{progress}$.
+$R_{hand\ dist}=e^{-||x_{palm}-x_{cup\ rim}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ rim}$ is the position of the cup rim.
+$R_{progress}=-||x_{cup}-x_{block}||$, where $x_{cup}$ is the position of the cup, and $x_{block}$ is the position of the block.
+  • Pour: $R_{task}=\beta_{1}R_{hand\ dist}+\beta_{2}R_{progress}+\beta_{3}R_{cup\ orient}$.
+$R_{hand\ dist}=e^{-||x_{palm}-x_{cup\ rim}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ rim}$ is the position of the cup rim, which is grasped by the facilitating hand.
+$R_{progress}=-||x_{cupd}-x_{cupf}||$, where $x_{cupf}$ is the position of the cup rim grasped by the facilitating hand, and $x_{cupd}$ is the position of the cup rim grasped by the dominant hand.
+$R_{cup\ orient}=\mathbf{z}_{cupf}\cdot\mathbf{z}_{world}$, where $\mathbf{z}_{cupf}$ is the z-axis unit vector of the cup grasped by the facilitating hand, and $\mathbf{z}_{world}$ is the z-axis unit vector of the world frame.
+  • Twist lid: $R_{task}=\beta_{1}R_{orient}+\beta_{2}R_{twist}+\beta_{3}R_{finger\ dist}+\beta_{4}R_{hand\ dist\ penalty}$.
+$R_{orient}=\mathbf{z}_{bottle}\cdot\mathbf{z}_{world}$, where $\mathbf{z}_{bottle}$ is the z-axis unit vector of the bottle grasped by the facilitating hand, and $\mathbf{z}_{world}$ is the z-axis unit vector of the world frame.
+$R_{twist}=\theta_{lid}^{t}-\theta_{lid}^{t-1}$, where $\theta_{lid}^{t}$ is the current bottle-lid revolute joint angle, and $\theta_{lid}^{t-1}$ is the previous bottle-lid revolute joint angle.
+And $R_{finger\ dist}$ is the finger contact reward we adopted from a previous lid twisting work [19].
+$R_{hand\ dist\ penalty}=-min((||x_{f}-x_{d}||-0.1),0.0)$, where $x_{f}$ and $x_{d}$ are the position of the facilitating hand palm and the dominant hand palm.
+Domain Randomization
+The domain randomization details are shown in Table. 6
+Table 6: Domain Randomization Setup.
+Object: Friction (only for twist lid task) [0.5, 1.5]
+Hand: Friction (only for twist lid task) [0.5, 1.5]
+Object Pos Observation Noise $+\mathcal{N}(0,0.02)$
+Hand Joint Observation Noise $+\mathcal{N}(0,0.2)$
+Hand Pos Observation Noise $+\mathcal{N}(0,0.02)$
+Hand Orientation Observation Noise $+\mathcal{N}(0,0.05)$
+Action Noise (only for twist lid task) $+\mathcal{N}(0,0.1)$